### v4

# Black Hole from Entropy Maximization

Abstract
One quantum characterization of a black hole motivated by (local) holography and thermodynamics is that it maximizes thermodynamic entropy for a given surface area. In the context of quantum gravity, this could be more fundamental than the classical characterization by a horizon. As a step, we explore this possibility by solving the 4D semi-classical Einstein equation with many matter fields, and find a picture of a black hole. For spherical static highly-excited configurations, we apply local typicality and estimate the entropy including self-gravity to derive its upper bound. The saturation condition uniquely determines the entropy-maximized configuration: self-gravitating quanta condensate into a radially-uniform dense configuration with no horizon, where the self-gravity and a large quantum pressure induced by the curvatures are balanced and no singularity appears. The interior metric is a self-consistent and non-perturbative solution for Planck’s constant. The maximum entropy, given by the volume integral of the entropy density, agrees with the Bekenstein-Hawking formula through self-gravity, deriving the Bousso bound for thermodynamic entropy. Finally, 10 future prospects are discussed, leading to the speculative view that the configuration represents semi-classically a quantum-gravitational condensate with holographic bulk dynamics.
I Introduction
What is a black hole in quantum theory? The answer is still unknown. Current observational data have not yet shown anything about the interior of a black hole or even confirmed the existence of a horizon LIGO ; EHT ; Cardoso . We have not yet found a theoretical description, fully consistent with quantum theory, that dynamically resolves both the information problem and the singularity, which should be rooted in the quantum nature of gravity. The geometrical characterization of black holes by their horizons was originally based on classical dynamics. However, now that the quantum properties of black holes Bekenstein ; Hawking have been discovered, it should be natural to consider that black holes are essentially quantum objects consisting of (still unknown) microscopic degrees of freedom, and that there is no a priori reason to follow the classical geometric definition in the context of quantum gravity, where spacetime fluctuates quantum mechanically. Therefore, a possible approach is to search for more appropriate and quantum definitions of black holes, to explore their identity, and to study the above problems and quantum gravity.
One quantum characterization of a black hole is that it maximizes thermodynamic entropy Dvali3 ; Oriti1 . This is not yet fully understood, but is motivated by a variety of facts. First, the microscopic origin of thermodynamic entropy is quantum: $S=\log\Omega$, where $\Omega$ is the number of quantum states $\{|\psi\rangle\}$ consistent with fixed macroscopic quantities. Therefore, the characterization by entropy is valid in a fully quantum context. Second, gravity is universal in that anything with energy attracts each other. In a strong gravity limit, any spherical configuration with mass $\frac{R}{2G}$ will collapse to a black hole with size $R$. This implies that a black hole can be considered as a macroscopic state with maximum entropy according to the second law of thermodynamics. Third, in the case of uncharged spherical symmetry, the Bekenstein-Hawking formula saturates the Bousso bound $S\leq\frac{{\cal A}}{4l_{p}^{2}}$ Bousso1 : a conjecture that the entropy $S$ inside a finite region is bounded by the boundary surface area ${\cal A}$, which has proposed holographic principle Bousso2 ($l_{p}\equiv\sqrt{\hbar G}$).111 Note here that the value of the Bekenstein-Hawking formula is determined by conserved charges such as ADM energy and does not depend on the presence of a horizon. See also Sec.VII. From these, the entropy maximization should be a candidate quantum characterization of black holes, eventually leading to a quantum description without classical geometry.
Then, what macroscopic quantities should we fix for entropy of a self-gravitating system? When calculating entropy of a non-gravitating object microcanonically, one fixes the energy and volume and counts up quantum states consistent with them. In a self-gravitating system, however, energy and volume are related to each other by the Einstein equation, and the two cannot be fixed independently. Rather, motivated by (local) holography Bousso2 ; Laurent , it should be natural to fix the surface area ${\cal A}$ of the boundary of a finite region, consider the entropy of the interior, and find the maximizing configuration Dvali3 ; Oriti1 .
[…]
Thus, as a first step toward a full quantum description, we consider spherically-symmetric static configurations and explore the characterization that a black hole maximizes thermodynamic entropy for a given surface area ${\cal A}=4\pi R^{2}$ $(R\gg l_{p})$, by using the 4D semi-classical Einstein equation, the self-consistent equation in a mean-field approximation of quantum gravity BD ; Kiefer :
$G_{\mu\nu}=8\pi G\langle\psi|T_{\mu\nu}|\psi\rangle,$ (1)
where gravity is described by classical metric $g_{\mu\nu}$ and matter by quantum operators. We then construct the interior metric $g_{\mu\nu}^{*}$ that maximizes the entropy, and find a candidate picture of quantum black holes. (Here, $X^{*}$ or $X_{*}$ denotes a quantity $X$ for the entropy-maximized configuration.) Note that we do not assume a priori that a black hole has a horizon or that the maximum entropy is given by the Bekenstein-Hawking formula.
In the framework of (1), a spherical static configuration for an excited state $|\psi\rangle$ can be considered as a collection of many excited quanta in $|\psi\rangle$ with self-gravity $g_{\mu\nu}$ satisfying (1) (see Fig.1).
Figure 1: A spherical static configuration with size $R$ as a collection of excited quanta in $(g_{\mu\nu},|\psi\rangle$) satisfying (1).
Then, we evaluate the thermodynamic entropy $S$ including the effect of self-gravity $g_{\mu\nu}$ by assuming the phenomenological form Francesco ; Zubarev ; FMW ; BFM :
[…]
where $\tilde{s}^{\mu}$ is the conserved entropy current, $\Sigma$ is chosen as a spacelike hypersurface orthogonal to the timelike Killing vector, and $s(r)$ is the entropy density per proper radial length. [^3] See (48) for the relation of $s$ and $\tilde{s}^{\mu}$.
Considering the pure-state formulation of (1) (and a future study of the unitary evolution in the information problem), it should be natural to evaluate the entropy density $s(r)$ by a pure-state method. Here, typicality Goldstein ; Popescu ; Sugita ; Reimann states that a subsystem typically behaves as a thermal state in a randomly-chosen state $|\psi\rangle$ from a sub-Hilbert space that is consistent with a given macroscopic parameter and has a sufficiently large density of states. We apply this idea to small subsystems compared to the radius of the curvatures, through the equivalence principle (Sec.II). For a typical and highly excited state $|\psi\rangle$, the energy density $\langle\psi|-T^{t}{}_{t}(r)|\psi\rangle$ behaves like a thermodynamic one, and the characteristic excitation energy of the typical quanta at $r$ corresponds to the local temperature $T_{loc}(r)$. This enables us to estimate the order of the magnitude of $s(r)$. Self-gravity is introduced through the Hamiltonian constraint, which relates $g_{\mu\nu}$ to $s(r)$. As a result, one metric $g_{\mu\nu}$ corresponds to a set of typical states $\{|\psi\rangle\}$ that have the same energy-momentum distribution, and the total entropy (2) is obtained as a functional of $g_{\mu\nu}$ for a given size $R$: $S=S[g_{\mu\nu};R)$. It can be seen that the entropy $S$ increases with the excitation energy at each point $r$. Here, in order for (1) to hold, we assume that the maximum excitation energy of a quantum is close to but smaller than the Planck energy.
The upper bound of the entropy then is derived from a semi-classical inequality required by the global static condition: the global existence of the timelike Killing vector (Sec.III). Note that static configurations do not have a trapped surface, since the Killing vector would be spacelike inside a trapped surface. Also, the bound leads to the Bekenstein bound Bek_bound including self-gravity.
Solving the saturation condition for the entropy bound under the maximum excitation and using the consistency with local thermodynamics, we find the entropy-maximized configuration uniquely (Sec.IV). The interior metric $g_{\mu\nu}^{*}$ is given by
$ds^{2}=-\frac{\sigma\eta^{2}}{2r^{2}}e^{-\frac{R^{2}-r^{2}}{2\sigma\eta}}dt^{2}+\frac{r^{2}}{2\sigma}dr^{2}+r^{2}d\Omega^{2},$ (3)
which is applied for $\sqrt{\sigma}\lesssim r\leq R$. Two parameters $\sigma=\mathcal{O}(nl_{p}^{2})$ and $\eta=\mathcal{O}(1)$ can be fixed by solving (1), where $n$ is a $\mathcal{O}(1)$ number to be large. [^4] Here, $\mathcal{O}(1)$ means $\mathcal{O}(r^{0})$ or $\mathcal{O}(R^{0})$ for $r,R\gg l_{p}$. We can check that for a theory with many matter fields, (3) is a non-perturbative and self-consistent solution of (1) in $\hbar$, leading to the species bound Dvali1 ; Dvali2 .
Geometrically, (3) is approximately a warped product of $AdS_{2}$ with radius $\sim\sqrt{n}l_{p}$ and $S^{2}$ with radius $r$. Physically, this shows that self-gravitating quanta with near-Planckian excitation energy condensate to form a radially-uniform dense configuration, as in Fig.2.
Figure 2: Black hole as the semi-classical gravity condensate with the maximum entropy $S_{max}=\frac{\mathcal{A}}{4l_{p}^{2}}$ and the interior metric (3): a collection of self-gravitating quanta with wavelength $\lambda_{*}(r)\sim\sqrt{n}l_{p}$ distributed uniformly in the proper radial length $\hat{r}$.
We call it a semi-classical gravity condensate. Here, the near-Planckian curvature inside induces quantum fluctuations of various modes, which generate a large tangential pressure $\langle T^{\theta}{}_{\theta}\rangle$ associated with the 4D Weyl anomaly, making the condensate locally anisotropic. This pressure supports the system against the strong self-gravity, exceeds the Buchdahl limit Buchdahl , and self-consistently keeps the curvature finite. As a result, the energy is distributed throughout the interior, and the small central region ($0\leq r\lesssim\sqrt{\sigma}$) beyond the description by (1) has only a small energy and can be assumed to be almost flat. Thus, no singularity appears.
The exterior part $(r\geq R)$ is described approximately by the Schwarzschild metric with ADM mass $\frac{a_{0}}{2G}(\gg m_{p}\equiv\sqrt{\hbar/G})$, which is related to the size $R$ as
$R=a_{0}+\frac{\sigma\eta^{2}}{2a_{0}}.$ (4)
This is close to but still outside $r=a_{0}$, and therefore, the configuration has no horizon but looks like a classical black hole from the outside. Indeed, it is almost black due to the exponentially large redshift of (3).
We can also obtain the gravity condensate by considering a formation process. Generically, for a given system, a configuration obtained by a reversible process should be thermodynamically typical and have the maximum entropy. In Refs.KMY ; KY1 ; KY4 ; KY5 , we considered a process in which thermal radiation comes together reversibly due to self-gravity in a heat bath at Hawking temperature, solved the self-consistent time evolution including the backreaction from Hawking-like radiation during the formation process, and obtained the metric (3). Therefore, the gravity condensate should be the most typical configuration with the maximum entropy according to the second law of thermodynamics, which is consistent with the above construction based on typicality.
Now, to evaluate explicitly $S_{max}\equiv S[g_{\mu\nu}^{*};R)$, we apply the Unruh effect (or the local temperature due to the particle creation inside) and thermodynamic relations locally to the interior metic (3) and obtain the entropy density $s_{*}(r)$ (Sec.V). We evaluate (2) and derive
$S\leq S_{max}=\frac{\mathcal{A}}{4l_{p}^{2}},$ (5)
where ${\cal A}\equiv 4\pi R^{2}=4\pi a_{0}^{2}+\mathcal{O}(1)$ for (4). Therefore, the maximum entropy $S_{max}$ coincides with the Bekenstein-Hawking formula. Here, the self-gravity changes the entropy (2) from the volume law to the area law Y1 .
This derives the Bousso bound for thermodynamic entropy (Sec.VI). Applying $\nabla_{\mu}\tilde{s}^{\mu}=0$ and a regularity condition to all the configurations in our class, we can show that (2) agrees with one evaluated on a light-sheet $\Sigma$ introduced in Ref.Bousso1 and thus is covariant. Therefore, (5) means that the gravity condensate is derived as the unique configuration saturating the Bousso bound. Furthermore, the interior metric (3) saturates the local sufficient conditions for the Bousso bound proposed in Refs.FMW ; BFM and thus has a holographic bulk dynamics.
We finally discuss 10 prospects for this picture of black holes (Sec.VII): role of self-gravity in holography, relation to other gravity-condensate models Dvali3 ; Oriti1 , gravity-condensate phase/state, path-integral evaluation of $S[g_{\mu\nu};R)$ BY1 ; BY2 ; Jacobson_vol , thermodynamic entropy vs entanglement entropy Minic ; Casini ; Jacobson_entangle , recovery of state-dependence in Hawking radiation KY2 , a picture of non-typical black holes, relation to the classical picture of black holes, gravitational field with finite entropy Jacobson ; Pad , and phenomenology echo ; CY . We then reach the speculative view that the gravity condensate represents semi-classically a mixture of gravity quanta and matter quanta.
II Estimation of entropy $S[g_{\mu\nu};R)$
II.1 Setup
We start with the setup. Suppose that an excited state $|\psi\rangle$ represents a spherically-symmetric static configuration of size $R$ such that (see Left of Fig.3)
$\langle\psi|T_{\mu\nu}(r)|\psi\rangle=\begin{cases}\neq 0&{\rm for}~{}~{}l_{p}\ll r\leq R\\ \approx 0&{\rm for}~{}~{}R\leq r.\end{cases}$ (6)
Here, we assume that $|\psi\rangle$ is excited enough to exceed possible negative energy contributions from vacuum fluctuations BD , making the total energy density positive ($\langle\psi|-T^{t}{}_{t}(r)|\psi\rangle>0$); we exclude a small center region where the semi-classical approximation may break down due to some quantum gravitational effect; and for simplicity, we consider the exterior part approximately vacuum, while a possible large backreaction effect from vacuum fluctuations around the Schwarzschild radius BD is taken into account in the interior part.
Then, we can set the metric by an ansatz:
$ds^{2}=\begin{cases}-\left(1-\frac{a(r)}{r}\right)e^{A(r)}dt^{2}+\left(1-\frac{a(r)}{r}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}{\rm for}~{}~{}l_{p}\ll r\leq R\\ -\left(1-\frac{a_{0}}{r}\right)dt^{2}+\left(1-\frac{a_{0}}{r}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}{\rm for}~{}~{}R\leq r.\end{cases}$ (7)
Here, $\frac{a(r)}{2G}$ is the Misner-Sharp mass inside $r$ Hayward , and $M_{0}\equiv\frac{a_{0}}{2G}\approx\frac{a(R)}{2G}$ is the ADM energy. [^5] The ADM energy $M_{0}$ is not fixed by hand but determined by the mass function $a(r)$ (together with an appropriate junction condition) for a given size $R$. Eventually, a set $(a(r),A(r),|\psi\rangle)$ is determined by solving the semi-classical Einstein equation (1) self-consistently. (See Appendix B for an example of the self-consistent analysis.) In the following, we suppose such a self-consistent configuration $(a(r),A(r),|\psi\rangle)$ satisfying (1).
We now consider this configuration from a microscopic point of view. It consists of a collection of many excited quanta in $|\psi\rangle$ as in Fig.1. Each excited quantum at $r$ may be in motion, but spherically and time averaged, it is stationary with respect to the timelike Killing vector $\partial_{t}$ and has the characteristic excitation energy, $\epsilon(r)(>0)$, measured locally.
II.2 Local typicality
According to the idea mentioned below (2), we estimate the entropy density $s(r)$ by using typicality locally. Let us first focus on a spherical subsystem at $r$ with width $\Delta\hat{r}\lesssim\mathcal{R}(r)^{-\frac{1}{2}}$ (see Fig.3).
Figure 3: Left: A self-consistent configuration $(g_{\mu\nu},|\psi\rangle)$. Right: A spherical subsystem with width $\Delta\hat{r}$ where typicality holds.
Here, $\mathcal{R}(r)$ denotes the order of the magnitude of the curvatures of the interior metric in (7), and $(\hat{t},\hat{r})$ is the local coordinate around $r$ with $d\hat{t}=\sqrt{-g_{tt}(r)}dt$ and $d\hat{r}=\sqrt{g_{rr}(r)}dr$. Therefore, the bulk of the subsystem can be considered locally flat. If $|\psi\rangle$ is sufficiently excited, a quantum around $r$ with $\epsilon(r)$ has a short wavelength such that
$\lambda(r)\sim\frac{\hbar}{\epsilon(r)}\lesssim\Delta\hat{r}\lesssim\mathcal{R}(r)^{-\frac{1}{2}}.$ (8)
Then, applying typicality Goldstein ; Popescu ; Sugita ; Reimann due to the equivalence principle, the spherical subsystem behaves like a local equilibrium in the radial direction such that the local energy $4\pi r^{2}\langle\psi|T^{\hat{t}\hat{t}}(r)|\psi\rangle\Delta\hat{r}$ agrees with the thermodynamic one:
$\langle\psi|-T^{t}{}_{t}(r)|\psi\rangle\approx\langle-T^{t}{}_{t}(r)\rangle_{th},$ (9)
666(9) is valid only for spherically symmetric averages. We do not consider only configurations that are in local equilibrium in the strict sense, such as those corresponding to local Gibbs states Zubarev , nor do we consider only static configurations where Tolman’s law holds rigorously Francesco . See also Sec.V.1.
and that $\epsilon(r)$ can be estimated by the local temperature $T_{loc}(r)$:
$\epsilon(r)\sim T_{loc}(r),$ (10)
since the local temperature governs the local energy scale for a highly excited state. [^7] We do not consider conserved charges involving chemical potentials. Here, we use $T^{\hat{t}\hat{t}}=-T^{\hat{t}}{}_{\hat{t}}=-T^{t}{}_{t}$ from $g^{\hat{t}\hat{t}}=-1$; $\langle-T^{t}{}_{t}(r)\rangle_{th}$ is the thermodynamic energy density given by a function of $T_{loc}(r)$ Landau_SM ; and $T_{loc}(r)$ is determined self-consistently by (1) (see Appendix A and Sec.V.1 for examples).
In general, the order of the magnitude of the entropy in a system with a small volume $\Delta V$ and a high temperature $T$ can be estimated as $\sim\langle-T^{t}{}_{t}\rangle_{th}\Delta V/T$. For example, in a ultra-relativistic spherical fluid with volume $4\pi r^{2}\Delta\hat{r}$ (including self-gravity), the Stefan-Boltzmann law holds approximately Landau_SM ; fluidbook , and the entropy is given by $s(r)\Delta\hat{r}=\frac{4\langle-T^{t}{}_{t}(r)\rangle_{th}}{3T_{loc}(r)}4\pi r^{2}\Delta\hat{r}$ (see (76)). Note here that the non-locality of entropy BFM ; BCFM is considered for the width (8). Therefore, using this, (9) and (10), the order of the magnitude of the entropy in our subsystem can be estimated by $s(r)\Delta\hat{r}\sim\frac{\langle\psi|-T^{t}{}_{t}(r)|\psi\rangle}{\epsilon(r)}4\pi r^{2}\Delta\hat{r}$, leading to [^8] One might think that in a pure state, entropy is zero. Microscopically, thermodynamic entropy is given by the number of possible states consistent with fixed macroscopic parameters (most such states are typical), which is state-independent Landau_SM . Physically, for example, a cold atomic system in a pure state will develop thermal behavior after a quench process, such that macroscopic quantities such as energy density have the same value as their thermal expectation values. Thus, assuming local typicality, the typical (i.e., thermal) behavior of the energy density can be used to estimate the entropy density in this phenomenological method. (See Sec.VII for microscopic methods.)
$s(r)\sim\frac{4\pi r^{2}\langle\psi|-T^{t}{}_{t}(r)|\psi\rangle}{\epsilon(r)}.$ (11)
For the above typicality-based evaluation to be valid, the subsystem must have a sufficiently large density of states, i.e. the entropy of the subsystem must be large:
$N(r)\equiv s(r)\Delta\hat{r}\gg 1.$ (12)
Here, using (11) and the total local energy of the subsystem $\Delta E_{loc}\equiv 4\pi r^{2}\langle T^{\hat{t}\hat{t}}(r)\rangle\Delta\hat{r}$, we have $N(r)\sim\frac{\Delta E_{loc}}{\epsilon(r)}$, which can be considered as the occupation number of excited quanta in the subsystem (see again Fig.3). Thus, the condition (12) should hold if we consider a highly excited state and a theory with many degrees of freedom. As we will see later, this is the case.
II.3 Self-gravity
We now introduce the effect of self-gravity. This is achieved by using the Hamiltonian constraint $\mathcal{H}=0$ ($G_{tt}=8\pi G\langle\psi|T_{tt}|\psi\rangle$) in the interior metric of (7):
$\partial_{r}a(r)=8\pi Gr^{2}\langle\psi|-T^{t}{}_{t}(r)|\psi\rangle.$ (13)
Applying this to (11), we have
$s(r)\sim\frac{\partial_{r}a(r)}{2G\epsilon(r)},$ (14)
which relates geometry and entropy. This leads to an interesting expression for the Misner-Sharp mass:
[…]
From this, the interior metric of (7) and (11), the total entropy (2) can be estimated as
[…]
which provides the entropy $S[g_{\mu\nu};R)$ for the self-consistent solution to (1). Thus, the Hamiltonian constraint (13) relates one metric $g_{\mu\nu}$ and a set of the typical states $\{|\psi\rangle\}$ that have the same thermodynamic energy density $\langle-T^{t}{}_{t}(r)\rangle_{th}$, and the finite entropy (16) is obtained. [^9] Here are some comments on the subtleties of the above entropy evaluation.We use only the geometrical static condition and local consistencies with thermodynamics, and do not assume a priori a global thermodynamic equilibrium. This is motivated by two facts. More physically, even for static configurations, one should consider their formation processes. Generically, mechanical and global/local thermodynamic equilibria are different due to differences in relaxation time scales Landau_SM . In particular, we are now considering self-gravity. Therefore, it is necessary to consider time-delay effects in the formation process and discuss in which equilibrium state the configuration is for the time scale under consideration. The other is that, a notion of global thermodynamic equilibrium in self-gravitating systems, leading often to thermodynamical instability from negative specific heat Antonov ; Lynden , is still controversial Landau_SM ; Pad_thermo , and in particular, one consistent with (1) is not known. Furthermore, we do not assume a condition of isotropic fluid Green ; Xia (like thermal radiation Sorkin ) because it is not clear a priori whether such fluid maximizes entropy including self-gravity dynamically. Thus, we have estimated $S[g_{\mu\nu};R)$, (16), by using only the geometric static condition and the local consistency with thermodynamics. For the case of $S_{max}$, we will check the validity of this treatment by constructing the self-consistent solution $(g_{\mu\nu}^{*},|\psi\rangle_{*})$ of (1) (Sec.IV.2). In Sec.VII we will discuss its global equilibrium briefly, and in Appendix D we will argue that Tolman’s law does not lead to maximum entropy.
The estimation (16) shows that for a given $s(r)$, the largest $\epsilon(r)$ at each $r$ leads to the largest $S$. More precisely, according to the second law within each spherical subsystem, the entropy density $s(r)$ should be an increasing function of the local temperature $T_{loc}(r)$ Landau_SM , and from (10), the maximum local temperature at each $r$ provides the maximum entropy $S_{max}$ for a given size $R$. This is consistent with ordinary thermodynamics without self-gravity Landau_SM , but it is a non-trivial result because self-gravity is included here.
Then, what is the maximum excitation energy? In order for the semi-classical description to be valid, the characteristic excitation energy $\epsilon(r)$ (or the local temperature $T_{loc}(r)$ from (10)) must satisfy Pad_Lim ; Caianiello ; Brandt
$\epsilon(r)\leq\epsilon_{max}\sim\frac{m_{p}}{\sqrt{n}}$ (17)
with $n=\mathcal{O}(1)\gg 1$, a large number to be determined.
III Upper bound
We derive the upper bound for $S[g_{\mu\nu};R)$. A static spacetime has a timelike Killing vector globally, indicating that there is no trapped surface Bousso2 ; No_trap . This condition can be expressed at a semi-classical level as
$\lambda(r)\lesssim\sqrt{g_{rr}(r)}(r-a(r))$ (18)
in the interior metric of (7): a quantum that constitutes a part at $r$ must be at least its wavelength out of the Schwarzschild radius $a(r)$ of the energy inside it Sorkin . (A dynamical origin of (18) will be discussed in Sec.IV.3.1.) Using (11) and $\lambda\sim\frac{\hbar}{\epsilon}$, we can rewrite (18) as
$s(r)\lesssim\frac{1}{\hbar}\sqrt{g_{rr}(r)}(r-a(r))4\pi r^{2}\langle\psi|-T^{t}{}_{t}(r)|\psi\rangle,$ (19)
whose meaning will be discussed in Sec.VI.2.1. From this, (2) and (7), we can calculate
$S$ $\lesssim\frac{4\pi}{\hbar}\int^{R}_{0}drr^{2}g_{rr}(r)(r-a(r))\langle\psi|-T^{t}{}_{t}(r)|\psi\rangle$
$=\frac{4\pi}{\hbar}\int^{R}_{0}drr^{3}\langle\psi|-T^{t}{}_{t}(r)|\psi\rangle.$ (20)
We apply (13) and obtain the upper bound:
[…]
which holds universally for any configuration in our class.
Here, we can derive from (III) the Bekenstein bound Bek_bound including self-gravity:
$S\lesssim\frac{RM_{0}}{\hbar},$ (22)
where we have used $4\pi\int^{R}_{0}drr^{3}\langle-T^{t}{}_{t}(r)\rangle\leq 4\pi R\int^{R}_{0}drr^{2}\langle-T^{t}{}_{t}(r)\rangle=R\frac{a(R)}{2G}\approx R\frac{a_{0}}{2G}\equiv RM_{0}$ from (13). [^10] In Ref.Sorkin , (III) and (22) were obtained only for thermal radiation, while our derivation can be applied to more general cases. Note that $G$ doesn’t appear here, but this is a result of the dynamics of gravity because we need $\mathcal{H}=0$ (13) to obtain the ADM energy $M_{0}$.
IV Entropy-maximized configuration
We take a strong-gravity limit in a consistent way and find the entropy-maximized configuration $g_{\mu\nu}^{*}$ that saturates the bound (21), to obtain a quantum picture of black holes. Here, (21) is just an order estimate, but remarkably, the saturation condition and the consistency with our arguments so far can determine the functional form of $g_{\mu\nu}^{*}(r)$ uniquely, except for two constant parameters. They can be fixed by solving (1) self-consistently.
IV.1 Saturating energy distribution $a_{*}(r)$
Let us first find the energy distribution $a_{*}(r)$ that saturates the inequality (21). Squaring the saturation condition for (18) and using (7) and $\lambda(r)\sim\frac{\hbar}{\epsilon(r)}$, we have
$\hbar^{2}\epsilon_{*}(r)^{-2}\sim r(r-a_{*}(r)).$ (23)
The estimation (16) and bound (17) mean that the maximum entropy is obtained by setting $\epsilon_{*}(r)=\epsilon_{max}$. Therefore, (23) becomes $nl_{p}^{2}\sim r(r-a_{*}(r))$, leading to
$a_{*}(r)=r-\frac{2\sigma}{r}$ (24)
with $\sigma=fnl_{p}^{2}$ ($f>0$: a constant of $\mathcal{O}(1)$). Thus, the maximum entropy $S_{max}\equiv S[g_{\mu\nu}^{*},R)$ is estimated from (21) and (24) as
[…]
where $a_{0}\approx a_{*}(R)\approx R$. (In Sec.V.2, the relation of $a_{0}$ and $R$ will be determined explicitly.) Note that the saturation of both (17) and (18) corresponds to the strong-gravity limit.
$a_{*}(r)$ represents a radially uniform configuration in that the entropy density is constant:
$s_{*}(r)\sim\frac{\sqrt{n}}{l_{p}},$ (26)
where we have applied $\epsilon(r)=\epsilon_{max}$ and (24) to (14). This means that $\mathcal{O}(\sqrt{n})$ bit of information is packed per the Planck length constantly KY2 ; Y1 . Thus, a necessary condition for the entropy maximization is the radial uniformity. Conversely, as a result of radial uniformity, (24) can be uniquely obtained as a solution satisfying the entropy maximization condition (23), without using $\epsilon(r)=\epsilon_{max}$ (see Appendix C). Therefore, from these two arguments, a necessary and sufficient condition for the entropy maximization is radial uniformity. This is a non-trivial result for a self-gravitating system.
We here examine in which region the energy distribution (24) is valid. As we will see in (35), the width $\Delta\hat{r}_{*}$ of the subsystem is almost the same as $\lambda_{*}(r)\sim\frac{\hbar}{\epsilon_{max}}\sim\sqrt{n}l_{p}$. From this, (26) and the assumption $n\gg 1$, then the number (12) of excited quanta in each subsystem is large indeed:
$N^{*}(r)\sim n\gg 1.$ (27)
On the other hand, using (13) and (24), we have the energy density $\langle-T^{t}{}_{t}\rangle_{*}\approx\frac{1}{8\pi Gr^{2}}$ and the local energy $\Delta E_{loc}^{*}(r)\equiv 4\pi r^{2}\langle-T^{t}{}_{t}\rangle_{*}\Delta\hat{r}_{*}\sim\sqrt{n}m_{p}$. Therefore, we must have at least $\frac{a_{*}(r)}{2G}\gtrsim\sqrt{n}m_{p}$, meaning that (24) holds only in
$\sqrt{n}l_{p}\lesssim r\leq R.$ (28)
For $n\gg 1$, this is consistent with the assumption we have made below (6).
IV.2 Determination of the interior metric $g_{\mu\nu}^{*}$
We determine the metric $g_{\mu\nu}^{*}$. First, (24) fixes $g_{rr}^{*}(r)=\frac{r^{2}}{2\sigma}$ in the interior metric of (7). Next, from (28) and the condition $n\gg 1$, we can focus on the asymptotic form $A(r)=Cr^{k}$ for $r\gg l_{p}$. Here, it should be natural to assume that there is no length scale in $g_{\mu\nu}^{*}$ except for $l_{p}$, since the saturating configuration consists of excited quanta with $\epsilon_{max}\sim\frac{m_{p}}{\sqrt{n}}$. Then, $C$ is a constant of $\mathcal{O}(1)$ which can depend on $l_{p}$ but not on $a_{0}$. Indeed, this will lead to a self-consistent solution of (1).
To find a physical value of $k$, we calculate
$-G^{t}{}_{t}$ $=\frac{1}{r^{2}}+\frac{2\sigma}{r^{4}},$ (29)
$G^{r}{}_{r}$ $=-\frac{1}{r^{2}}-\frac{2\sigma}{r^{4}}+2Ck\sigma r^{k-4},$ (30)
$G^{\theta}{}_{\theta}$ $=\frac{2\sigma}{r^{4}}+C(k-3)k\sigma r^{k-4}+\frac{1}{2}C^{2}k^{2}\sigma r^{2k-4},$ (31)
and use two self-consistencies: thermodynamics and semi-classicality. Because we have considered quanta consistent with thermodynamics, the pressures must be positive Landau_SM . This and (30) require $C>0$ and $k-4\geq-2$, that is, $k\geq 2$. Also, we are focusing only on semi-classical configurations satisfying (17), which means that the curvatures $\mathcal{R}(r)$ must be at most of $\mathcal{O}(r^{0})$ for $r\gg l_{p}$. This restricts the highest term, the third one of (31), such that $2k-4\leq 0$, that is, $k\leq 2$. Therefore, we can only have $k=2$. From dimensional analysis, we can set $C=\frac{1}{2\sigma\eta}$ with a $\mathcal{O}(1)$ dimensionless constant $\eta>0$.
Thus, we have reached uniquely the interior metric $g_{\mu\nu}^{*}$ with the maximum entropy in the class of spherically-symmetric static spacetime:
$ds^{2}=-\frac{2\sigma}{r^{2}}e^{\frac{r^{2}}{2\sigma\eta}+A_{0}}dt^{2}+\frac{r^{2}}{2\sigma}dr^{2}+r^{2}d\Omega^{2},$ (32)
which is valid for the range (28). $A_{0}$ is a constant for connection to the exterior metric (see (46)).
Note that (32) can be obtained in many ways: adiabatic formation in a heat bath of Hawking temperature KMY ; KY1 ; KY4 ; KY5 , consistency with the Bekenstein-Hawking formula Y1 , and typical configuration in a solution space of (1) HKLY . In this sense, (32) is robust, and the present argument provides a derivation that characterizes it as the entropy-maximizing configuration.
We now check that the metric (32) satisfies (1) self-consistently. We here give an outline of the proof. (See Appendix B for a short review and Ref.KY4 for the details.) We first consider, say, $n_{s}(\gg 1)$ scalar fields in the background metric (32) and solve the matter field equations with a perturbative technique by employing the fact that the metric is a warped product of $AdS_{2}$ with radius $L\equiv\sqrt{2\sigma\eta^{2}}$ and $S^{2}$ with radius $r$ (see the Ricci scalar in (34)). [^11] More precisely, we have $R_{RicciScalar}=-\frac{2}{L^{2}}+\frac{2}{r^{2}}$, and we can check that the metric (32) is equivalent locally to $ds^{2}=\frac{L^{2}}{z^{2}}(-dt^{2}+dz^{2})+r(z)^{2}d\Omega^{2}$ KY4 . We then use dimensional regularization to evaluate the renormalized energy-momentum tensor $\langle\psi|T_{\mu\nu}|\psi\rangle_{*}$ for a typical state $|\psi\rangle_{*}$ in which only s-waves are excited with $\sim\epsilon_{max}$ and the other modes are in the ground state $|0\rangle$ of (32). [^12] Note that $|\psi\rangle_{*}$ is not the local Gibbs state Zubarev with $T_{loc}\sim\epsilon_{max}$. We finally compare both sides of (1) to find the self-consistent values:
$\sigma=\frac{n_{s}l_{p}^{2}}{120\pi\eta^{2}},~{}~{}1\leq\eta<2.$ (33)
We can check that (33) is consistent with 4D Weyl anomaly BD (see Appendix B). Thus, we conclude that (32) with (33) is the non-perturbative solution of (1) with $n_{s}$ scalar fields in the sense that the limit $\hbar\to 0$ cannot be taken in (32) and (34).
We here discuss two points about the solution.
IV.2.1 Meaning of $n$ and species bound
We discuss the meaning of $n$. Comparing $\sigma=fnl_{p}^{2}$ to (33), we have $n=n_{s}$, the number of scalar fields. In the case where matter fields are conformal, we can determine the self-consistent value of $\sigma$, which is different from (33) and shows $n=c_{W}$ KY1 ; KY3 ; KY5 . Here, $c_{W}$ is the coefficient of the square of Weyl tensors in the 4D Weyl anomaly BD . Therefore, $n$ represents the number of the degrees of freedom in the theory that can contribute to entropy (see (26) and (41)). If we consider a theory with many species of fields, the condition $n\gg 1$ is satisfied. On the other hand, $n$ has been introduced in (17) as a parameter characterizing the maximum energy $\epsilon_{max}$ for which a semi-classical description is valid. Thus, (17) agrees with the species bound Dvali1 . This is the result of solving (1) non-perturbatively, which is another derivation of the species bound. Note that the species bound also appears from the validity of the Bekenstein bound Dvali2 , while we first derive the upper bound (21) and then reach the species bound. This may imply that there is an intrinsic relationship between the species bound and the entropy bound.
IV.2.2 No singularity
We can expect that there is no singularity. First, note that the interior metric (32) is valid only in the range (28). In a large $n$, we have the leading terms of the curvatures for $r\gg l_{p}$:
$R_{RicciScalar}=-\frac{2}{L^{2}},~{}~{}R_{\mu\nu}R^{\mu\nu}=\frac{2}{L^{4}},~{}~{}R_{\mu\nu\alpha\beta}R^{\mu\nu\alpha\beta}=\frac{4}{L^{4}},$ (34)
where $L\equiv\sqrt{2\sigma\eta^{2}}\sim\sqrt{n}l_{p}\gg l_{p}$. These are close to but still smaller than the Planck scale: $\mathcal{R}_{*}(r)\sim\frac{1}{nl_{p}^{2}}$. On the other hand, we have $\lambda_{*}(r)\sim\frac{\hbar}{\epsilon_{max}}\sim\sqrt{n}l_{p}$. Therefore, the width of the subsystems is minimum:
$\Delta\hat{r}_{*}\sim\lambda_{*}(r)\sim\mathcal{R}_{*}(r)^{-\frac{1}{2}}\sim\sqrt{n}l_{p},$ (35)
which satisfies the condition (8), albeit barely.
We next examine the small center region $0\leq r\lesssim\sqrt{n}l_{p}$, which cannot be described by the metric (32). The energy inside is estimated through (24) as $\frac{a_{*}(\sqrt{n}l_{p})}{2G}\sim\sqrt{n}m_{p}$, which is much smaller than $\frac{a_{0}}{2G}$, one expected in a classical case. Therefore, the center region cannot have a singularity appearing in the classical cases, and rather it should correspond to a small excitation of quantum-gravitational degrees of freedom like string (which can be confirmed only by a future development). In this paper, we assume that the center part is almost flat Y1 . [^13] The metric (32) becomes flat around $r\approx\sqrt{2\sigma}$: $ds^{2}|_{r\approx\sqrt{2\sigma}}\approx-e^{\frac{1}{\eta}+A_{0}}dt^{2}+dr^{2}+r^{2}d\Omega^{2}$, which is a flat metric by redefining time coordinate $t$. Also, we can consider the time evolution from formation to evaporation and see that the center region is kept flat, except for the final stage of the evaporation KY4 .
Thus, the non-perturbative solution (32) with (33) has no singularity. We will see in Sec.IV.3.1 that a quantum repulsive force is generated inside and plays a key role in the resolution of the singularity.
IV.3 Semi-classical gravity condensate
We study the configuration of (32). (See Ref.KY2 for aspects not discussed here.) First, it has through (1)
$\langle-T^{t}{}_{t}(r)\rangle_{*}=\frac{1}{8\pi Gr^{2}},$ $~{}~{}\langle T^{r}{}_{r}(r)\rangle_{*}=\frac{2-\eta}{\eta}\langle-T^{t}{}_{t}(r)\rangle_{*},$
$\langle T^{\theta}{}_{\theta}(r)\rangle_{*}$ $=\frac{1}{16\pi G\eta^{2}\sigma},$ (36)
[…]
Noting the radial uniformity (discussed in Sec.IV.1), we thus reach the picture shown in Fig.2. The metric (32) represents a self-gravitating condensate consisting of the excited quanta with $\epsilon_{max}$, distributed uniformly in $\hat{r}$ direction, and the vacuum fluctuations (semi-classical gravity condensate). It is dense in the sense that it has the large pressure and curvatures. The exterior geometry ($r\geq R$) is approximated by the Schwarzschild metric in (7), where the curvature is small. We can check here that the curvatures jump at the surface (located at (45)) in a mild manner to keep as much the interior uniformity as possible, consistent with Israel’s junction condition Poisson (see Sec.7 in Ref.Y1 for the details).
Let us now find a rough form of $A_{0}$ and consider what it means. First, we note from (24) that $a_{0}\approx a_{*}(R)=R-\frac{2\sigma}{R}$, indicating that the surface exists at $r=R=a_{0}+\mathcal{O}\left(\frac{nl_{p}^{2}}{a_{0}}\right)$ (see (45) for the precise one). For the interior metric (32) to connect at $r=R$ to the exterior Schwarzschild metric in (7), the induced metric on $r=R$ must be continuous Poisson , which requires $-g_{tt}(R)=1-\frac{a_{0}}{R}\approx\frac{\mathcal{O}(nl_{p}^{2})}{R^{2}}=\frac{2\sigma}{R^{2}}e^{\frac{R^{2}}{2\sigma\eta}+A_{0}}$. Then, we obtain
$g_{tt}^{*}(r)=-\frac{2K\sigma}{r^{2}}e^{-\frac{R^{2}-r^{2}}{2\sigma\eta}},$ (37)
where $K$ is a $\mathcal{O}(1)$ number to be fixed. This is an exponentially large redshift KMY . To see this effect, setting $r=R-\Delta r$ ($\Delta r\ll R$), (37) becomes $-g_{tt}^{*}(r)\sim e^{-\frac{R\Delta r}{\sigma\eta}}$. This indicates that due to the strong redshift, from the outside, a deep region with $\Delta r\gg\mathcal{O}\left(\frac{nl_{p}^{2}}{a_{0}}\right)$ is almost frozen in time. As a result, the information carried by the excited quanta is kept inside for an exponentially long time KY2 , and the gravity condensate looks almost black. (We will discuss its phenomenological effect in Sec.VII.)
IV.3.1 Formation process
We discuss how the gravity condensate is formed by a physical process. See Fig.4.
Figure 4: Adiabatic formation of the semi-classical gravity condensate in a heat bath of Hawking temperature.
Suppose that a collection of many spherical radiations at Hawking temperature comes together slowly due to self-gravity. This corresponds to the adiabatic formation process in a heat bath and can be done operationally by slowly changing the temperature of the heat bath to match the Hawking temperature of the energy at each stage KY1 .
We can use (1) and make a self-consistent model that describes this process KMY ; KY1 ; KY4 ; KY5 ; Y1 . We first note that in this formation process, the metric changes in time, and that generically, particle creation occurs in a time-dependent spacetime BD ; Barcelo . Then, we can solve approximately quantum fields propagating in this metric to show that particles at the Hawking temperature are created in each part of the radiation. Considering the time evolution of both the collapsing matter and the backreaction from the particle creation by (1), [^15] It is often believed that the backreaction from evaporation is negligible when considering the formation process of a black hole with mass $\frac{a_{0}}{2G}$ since the time scale of the evaporation $\Delta t\sim\frac{a_{0}^{3}}{l_{p}^{2}}$ is much longer than the time scale of the collapse $\Delta\tau\sim a_{0}$. However, this naive idea is wrong because $t$ and $\tau$ are the time coordinate at infinity and that of the comoving observer along the collapsing matter, respectively, and it does not make sense to compare such two time scales. In fact, one can examine the time evolution of the evaporation and collapse in a common time coordinate, and check that the both effects compete near $r=a_{0}$ and the dense structure is formed KMY ; KY4 . the gravity condensate will be formed with the maximum entropy (25). This result is consistent with the second law of thermodynamics because reversible processes generally lead to most typical configurations in thermodynamics.
In the heat bath with the fixed Hawking temperature $T_{H}=\frac{\hbar}{4\pi a_{0}}$, the incoming radiation from the bath and the outgoing one from the condensate are balanced, and the system stays stationary KY2 . When taken out of the bath into a vacuum, the condensate evaporates in a time scale $\Delta t\sim\frac{a_{0}^{3}}{\sigma}$ KMY ; KY1 .
In this formation process, we can see three interesting points. First, we can find the origin of the large tangential pressure in (IV.3). Let us imagine that self-gravity brings together a large number of excited modes representing radiation from the bath, and the condensate is gradually formed. As the curvature increases, vacuum fluctuations of other modes with various angular momenta are induced causing the pressure as a non-perturbative effect KY4 (see Appendix B for a review). This can also be understood through the 4D Weyl anomaly KY3 . Because of the pressure, the excited quanta are not concentrated in the center but are distributed throughout the interior. As a result, the curvature remains finite as in (34), the center has only small energy, and no singularity appears.
Second, the adiabatic formation is consistent with Bekenstein’s idea Bekenstein . At each stage where the mass is $\frac{a_{0}^{\prime}}{2G}$, a photon emitted from the bath of $T_{H}^{\prime}=\frac{\hbar}{4\pi a_{0}^{\prime}}$ comes to the gravity condensate of size $R^{\prime}=a_{0}^{\prime}+\mathcal{O}\left(\frac{nl_{p}^{2}}{a_{0}^{\prime}}\right)$. Since the wavelength $\sim a_{0}^{\prime}$ and the size $R^{\prime}$ are almost the same, the probability for the photon to enter the dense region is roughly one-half, leading to 1 bit of information Bekenstein . On the other hand, the photon approaches to $r=R^{\prime}$  [^16] The model in Ref.KMY ; KY4 shows explicitly that, as a result of the self-consistent time-evolution considering the point of footnote 15, the photon will naturally approach $r=a^{\prime}_{0}+\frac{2\sigma}{a^{\prime}_{0}}$ with emitting particles at temperature $T_{H}^{\prime}=\frac{\hbar}{4\pi a_{0}^{\prime}}$., and the blueshifted temperature becomes $T_{loc}(r=R^{\prime})=\frac{T_{H}^{\prime}}{\sqrt{1-\frac{a_{0}^{\prime}}{R^{\prime}}}}\sim\frac{m_{p}}{\sqrt{n}}\sim\epsilon_{max}$. Therefore, this photon has $\epsilon_{max}$ and the 1 bit of information about its existence. Repeating this process, the uniform dense structure like Fig.2 appears, and the entropy follows the area law (25). In this sense, Bekenstein’s idea is realized, including the interior description at the level of (1).
Third, the semi-classical condition (18) is satisfied as a result of the time evolution by (1). As shown above, each photon approaches $r=R^{\prime}=a_{0}^{\prime}+\mathcal{O}\left(\frac{nl_{p}^{2}}{a_{0}^{\prime}}\right)$, which makes the right hand side of (18) into $\sqrt{g_{rr}(R^{\prime})}(R^{\prime}-a_{0}^{\prime})\sim\sqrt{n}l_{p}$. On the other hand, the photon has the energy $\epsilon_{max}$, which means that the left hand side of (18) becomes $\frac{\hbar}{\epsilon_{max}}\sim\sqrt{n}l_{p}$. Then, the condition (18) is saturated at each stage of the formation. Note here that the condition (18) leads to the entropy bound (21), and that the formation process is reversible thermodynamically. Thus, these may imply that the condition (18), and hence the bound (21), is a consequence of the second law of thermodynamics at the level of the semi-classical Einstein equation (1).
V Maximum Entropy $S_{max}$
Following the derivation in Ref.Y1 , we show that the maximum entropy $S_{max}\equiv S[g_{\mu\nu}^{*};R)$ agrees with the Bekenstein-Hawking formula including the coefficient $1/4$ for any type of matter fields, and then identify the relation of $R$ and $a_{0}$ from a thermodynamical argument.
V.1 Bekenstein-Hawking formula
V.1.1 Local temperature
We first find the local temperature $T_{loc}^{*}(r)$. We note that the excited quanta composing the part around $r$ of the gravity condensate with the metric (32) are in radially accelerated motion against gravity to stay there as in Fig.2. The required acceleration is
$\alpha_{*}(r)=\frac{1}{\sqrt{2\eta^{2}\sigma}}+\mathcal{O}(r^{-1}).$ (38)
Here, we have applied the metric (32) and the formula of the proper acceleration in the interior metric of (7): $\alpha(r)\equiv|g_{\mu\nu}\alpha^{\mu}\alpha^{\nu}|^{\frac{1}{2}}=\frac{\partial_{r}\log\sqrt{-g_{tt}(r)}}{\sqrt{g_{rr}(r)}}$, $\alpha^{\mu}\equiv u^{\nu}\nabla_{\nu}u^{\mu}$, and $u^{\mu}\partial_{\mu}=(-g_{tt}(r))^{-\frac{1}{2}}\partial_{t}$. Then, we have $\alpha_{*}(r)^{-1}\sim\sqrt{n}l_{p}\sim\mathcal{R}_{*}(r)^{-1/2}$ (from (35)), and the quanta can be considered accelerating in a locally flat subsystem. Therefore, we can apply the Unruh effect locally Jacobson ; Pad :
$T_{loc}^{*}(r)=\frac{\hbar\alpha_{*}(r)}{2\pi}=\frac{\hbar}{2\pi\sqrt{2\sigma\eta^{2}}}+\mathcal{O}(r^{-1}),$ (39)
which is $\sim\epsilon_{max}$ consistent with (10). Thus, the excited quanta behave typically like a local thermal state in the radial direction at the temperature (39). This result is kinematical and can be applied to any type of local degrees of freedom in the metric (32), due to the universality of Unruh effect. On the other hand, we have seen in Sec.IV.3.1 that (39) is obtained dynamically through the particle creation during the formation process (see Y1 for an explicit proof). Therefore, the kinematical and dynamical results coincide, and (39) is robust.
At first glance, the fact that $T_{loc}^{*}(r)$ is constant might appear to contradict Tolman’s law, while its naive application to (37) for $r\ll R$ would lead to $T_{loc}(r)=\frac{T_{H}}{\sqrt{-g_{tt}^{*}(r)}}\sim\frac{\hbar}{\sqrt{\sigma}}e^{\frac{R^{2}}{4\sigma\eta}}\gg\frac{m_{p}}{\sqrt{n}}$, violating the semi-classical condition (17). In general, Tolman’s law holds only if, in a stationary spacetime, thermal radiation (more generally, energy flow) can propagate between objects that are stationary with respect to each other within the considered time, including (if any) the effects of interaction with other modes and scattering by potentials Tolman ; Landau_SM . On the other hand, using the metric (32) with (37), the radial null geodesic equation is given by $\frac{dr(t)}{dt}=\pm\sqrt{K}\frac{2\sigma}{r^{2}}e^{-\frac{R^{2}-r^{2}}{4\sigma\eta}}$, which means that in the absence of interaction, it takes time $\Delta t\sim e^{a_{0}^{2}/l_{p}^{2}}$ for thermal radiation to travel a distance $\Delta r=\mathcal{O}(a_{0})$ inside.
Therefore, we can understand (39) as follows. Each part of the interior remains, due to the large redshift, at the temperature $T_{loc}(r)\sim\epsilon_{max}$ at the time of formation (see Sec.IV.3.1), and only during $\Delta t<\mathcal{O}(e^{a_{0}^{2}/l_{p}^{2}})$ can the gravity condensate exist consistent with Tolman’s law. In this sense, the entropy-maximizing configuration is not in global equilibrium but in radially local one  [^17] This non-global equilibrium makes the difference from the result of Ref.Oppenheim , and is consistent with the self-consistent state $|\psi\rangle_{*}$ (mentioned above (33)). See Sec.VII for discussions about a self-gravitating equilibrium state.. This is consistent with the result that Tolman’s law does not lead to maximum entropy (see Appendix D). Note that the time scale $\sim e^{a_{0}^{2}/l_{p}^{2}}$ is much longer than the evaporation time scale $\sim\frac{a_{0}^{3}}{l_{p}^{2}}$, and we can discuss the configuration physically.
V.1.2 Derivation of the entropy-area law
Now, in this local equilibrium, the 1D Gibbs relation
$T_{loc}s=\rho_{1d}+p_{1d}$ (40)
holds for $\rho_{1d}=4\pi r^{2}\langle-T^{\hat{t}}{}_{\hat{t}}\rangle$ and $p_{1d}=4\pi r^{2}\langle T^{\hat{r}}{}_{\hat{r}}\rangle$, since the configuration is uniform in the proper radial length $\hat{r}$ Groot . Also, from (IV.3), $p_{1d}^{*}=\frac{2-\eta}{\eta}\rho_{1d}^{*}$ plays the role of the equation of state because $\langle T^{\theta}{}_{\theta}\rangle_{*}$ comes from the vacuum and thus has no thermodynamic contribution. Indeed, this treatment is consistent with the Bousso bound (see Sec.VI.2.2). These together with (39) and (IV.3) provide
$s_{*}(r)=\frac{\rho_{1d}^{*}+p_{1d}^{*}}{T_{loc}^{*}}=\frac{1}{T_{loc}^{*}}\frac{2}{\eta}\rho_{1d}^{*}=\frac{2\pi\sqrt{2\sigma}}{l_{p}^{2}},$ (41)
which is $\sim\frac{\sqrt{n}}{l_{p}}$ and consistent with (26). Applying this and (32) to (2), we obtain the Bekenstein-Hawking formula Bekenstein ; Hawking :
[…]
where ${\cal A}\equiv 4\pi R^{2}=4\pi a_{0}^{2}+\mathcal{O}(1)$. This gives the precise version of (25), and therefore (21) and (42) mean (5).
Note that while the entropy density (41) depends on the species of fields and the equation of state through $\sigma$ in (33), $\sigma$ cancels out in (42), leading to the coefficient $1/4$. This is the result of the self-consistent self-gravity (32) and the universal temperature (39) Y1 .
V.2 Surface
We here determine the relation of the size $R$ and the total energy $M_{0}=\frac{a_{0}}{2G}$ from thermodynamics Y1 . Imagine that the gravity condensate is in equilibrium with a heat bath of temperature $T_{0}$ during $\Delta t<\mathcal{O}(e^{a_{0}^{2}/l_{p}^{2}})$ (see Right of Fig.4). From thermodynamic relation $T_{0}dS=dM_{0}$ and (42), the equilibrium temperature is determined as the Hawking temperature GH :
$T_{0}=\frac{\hbar}{4\pi a_{0}}.$ (43)
Now, radiation emitted from the bath at $r\gg a_{0}$ comes close to the surface at $r=R$, and according to Tolman’s law in the Schwarzschild metric, the blueshifted temperature is given by $\frac{T_{0}}{\sqrt{1-\frac{a_{0}}{R}}}$. On the other hand, the argument based on the internal structure determines the local temperature at $r\leq R$ as (39). Then, considering the interior and exterior parts as two thermodynamic phases, thermodynamics requires that the local temperature be continuous at the boundary, $r=R$ Landau_SM : [^18] Here, the energy flow is balanced between the condensate and the heat bath, and there is no net energy flow. Therefore, the latent heat at the boundary should be zero, and the local temperature should be continuous Nakagawa-Sasa .
$T_{loc}^{*}(R)=\frac{T_{0}}{\sqrt{1-\frac{a_{0}}{R}}}.$ (44)
This determines the relation of $R$ and $a_{0}$. Setting $R=a_{0}+\Delta$ ($\Delta\ll a_{0}$) and using (39) and (43), (44) becomes $\frac{\hbar}{2\pi\sqrt{2\sigma\eta^{2}}}=\frac{\hbar}{4\pi a_{0}}\frac{1}{\sqrt{\frac{\Delta}{R}}}\approx\frac{\hbar}{4\pi\sqrt{a_{0}\Delta}}$, leading to $\Delta=\frac{\sigma\eta^{2}}{2a_{0}}$. We thus obtain (4):
$R=a_{0}+\frac{\sigma\eta^{2}}{2a_{0}}.$ (45)
This satisfies Israel’s condition and can also be obtained from a mechanical argument Y1 . Note that the proper length $\sqrt{g_{rr}^{*}(R)}\frac{\sigma\eta^{2}}{2a_{0}}\sim\sqrt{n}l_{p}~{}(\gg l_{p})$ is independent of $a_{0}$.
Thus, using (45) and fixing $A_{0}$ with the same procedure to get (37), we obtain the complete form of $g_{\mu\nu}^{*}$:
$ds^{2}=-\frac{\sigma\eta^{2}}{2r^{2}}e^{-\frac{R^{2}-r^{2}}{2\sigma\eta}}dt^{2}+\frac{r^{2}}{2\sigma}dr^{2}+r^{2}d\Omega^{2},$ (46)
which is (3).
VI Bousso bound
In this section, we use the results obtained so far to derive the Bousso bound. We then confirm that the gravity condensate saturates not only the Bousso bound but also the local sufficient conditions proposed in the literature.
VI.1 Derivation of the Bousso bound
We show that the Bousso bound holds in our class of configurations. We first check that the entropy evaluated on a $t$-constant spacelike hypersurface $\Sigma_{t}$ inside the surface at $r=R$, which we have considered so far, agrees with one evaluated on a spherical ingoing null hypersurface $\Sigma_{L}$ starting from the surface at $r=R$ and converging at $r=0$ (a light sheet) Bousso1 . We suppose here that all configurations satisfy (1) and are regular. Applying $\nabla_{\mu}\tilde{s}^{\mu}=0$ to a finite spacetime region $V$ enclosed with $\Sigma_{t}$ and $\Sigma_{L}$ ($\partial V=\Sigma_{L}-\Sigma_{t}$ as in Fig.5)
Figure 5: A spacetime region $V$ enclosed with $\Sigma_{t}$ and $\Sigma_{L}$.
and using Gauss’ theorem, we obtain
[…]
Therefore, the entropy $S[g_{\mu\nu},R)$, (16), agrees with the covariant entropy. Then, combining this and (5), we obtain the Bousso bound for thermodynamic entropy.
Thus, from this result and the discussion so far, we conclude that in the class of spherically-symmetric static configurations, only the solution metric (46) saturates the Bousso bound.
VI.2 Local sufficient conditions for the Bousso bound
We study the relation of our discussion and the sufficient conditions proposed in Refs.FMW ; BFM for the Bousso bound and check the consistency.
First, we prepare ingredients to be used below. In a static spherical configuration with metric (7), the relation of entropy density $s$ and entropy current $\tilde{s}^{\mu}$ is as follows. For a $t$-constant hypersurface, we have $d\Sigma_{\mu}=-u_{\mu}\sqrt{g_{rr}}r^{2}\sin\theta drd\theta d\phi$ with $u_{\mu}dx^{\mu}=-\sqrt{-g_{tt}}dt$ Poisson . Then, (2) gives $s=4\pi r^{2}(-u_{\mu}\tilde{s}^{\mu})$, i.e.
$s(r)=4\pi r^{2}\sqrt{-g_{tt}(r)}\tilde{s}^{t}(r),$ (48)
and the entropy current is given by $\tilde{s}^{\mu}\partial_{\mu}=\tilde{s}^{t}\partial_{t}$.
We consider a spherical light sheet. A radially-ingoing null vector $k^{\mu}$ tangent to a geodesic generating it satisfies $0=g_{\mu\nu}k^{\mu}k^{\nu}=g_{tt}(k^{t})^{2}+g_{rr}(k^{r})^{2}$, leading to
$\left(\frac{k^{t}}{k^{r}}\right)^{2}=\frac{g_{rr}}{-g_{tt}}.$ (49)
Using this, we can calculate for a spherical static energy-momentum tensor $\langle T_{\mu\nu}(r)\rangle$
$\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu}$ $=\langle T_{tt}\rangle(k^{t})^{2}+\langle T_{rr}\rangle(k^{r})^{2}$
$=g_{rr}(k^{r})^{2}\left(\left(\frac{k^{t}}{k^{r}}\right)^{2}\frac{-g_{tt}}{g_{rr}}\langle-T^{t}{}_{t}\rangle+\langle T^{r}{}_{r}\rangle\right)$
$=g_{rr}|k^{r}|^{2}\left(\langle-T^{t}{}_{t}\rangle+\langle T^{r}{}_{r}\rangle\right).$ (50)
VI.2.1 Meaning of the upper bound (19)
A local sufficient condition, Eq.(1.9) in Ref.FMW (or one in Ref.Pesci ), is given by
$|-\tilde{s}^{\mu}k_{\mu}|\lesssim\frac{1}{\hbar}\Delta\zeta\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu},$ (51)
where $\Delta\zeta$ is the (finite) affine length of a geodesic along $k^{\mu}$ generating a light sheet in which quanta contributing to the entropy are included completely. (51) is a kind of light ray equivalent of Bekenstein’s bound (22), since $\Delta\zeta$ and $\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu}$ correspond to $R$ and $M_{0}$, respectively.
We here show that the static condition (19) (leading to the bound (21)) is a static spherical version of (51). First, we apply (48) to (19) and get
$\tilde{s}^{t}\lesssim\frac{1}{\hbar}\sqrt{\frac{g_{rr}}{-g_{tt}}}(r-a(r))\langle-T^{t}{}_{t}\rangle.$ (52)
We here have $|-\tilde{s}^{\mu}k_{\mu}|=|-\tilde{s}^{t}k_{t}|=\tilde{s}^{t}(-g_{tt})|k^{t}|$. Also, for excitations consistent with thermodynamics, $\langle-T^{t}{}_{t}\rangle\sim\langle T^{r}{}_{r}\rangle$ holds Landau_SM , and (VI.2) means $\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu}\sim g_{rr}|k^{r}|^{2}\langle-T^{t}{}_{t}\rangle$. Using these, (52) leads to
$|-\tilde{s}^{\mu}k_{\mu}|$ $\lesssim\frac{1}{\hbar}\sqrt{\frac{-g_{tt}}{g_{rr}}}\frac{|k^{t}|}{|k^{r}|}\frac{r-a(r)}{|k^{r}|}\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu}$
$=\frac{1}{\hbar}\frac{r-a(r)}{|k^{r}|}\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu},$ (53)
where (49) is applied. In terms of the affine parameter $\zeta$ of the geodesic, we have $k^{r}=\frac{dr(\zeta)}{d\zeta}$ and express $\frac{r-a(r)}{|k^{r}|}=\left|\frac{d\zeta}{dr}\right|(r-a(r))=\Delta\zeta$. Thus, (VI.2.1) provides (51) at each $r$ along the geodesic.
We discuss the meaning of this $\Delta\zeta$. In general, when considering a light sheet going to a horizon, the light sheet must not be continued to the interior, since the horizon entropy contains the contribution from objects that fell inside: Otherwise, it would be counted twice Bousso2 . In the geometry (7) without trapped surface, the “would-be horizon” for a point $r$ exists at $r=a(r)$. Therefore, our $\Delta\zeta$ represents the affine distance from a point to the “would-be end point”, and hence our $\Delta\zeta$ corresponds to that of Eq.(1.9) in Ref.FMW .
Thus, we conclude from the arguments so far that the spacetime $g_{\mu\nu}^{*}$ (46) is derived as a necessary condition for saturating (51).
VI.2.2 Exact saturation of a local sufficient condition by $g_{\mu\nu}^{*}$
We study another local sufficient condition, Eq.(3.5) in Ref.BFM :
$|k^{\mu}k^{\nu}\nabla_{\mu}\tilde{s}_{\nu}|\leq\frac{2\pi}{\hbar}\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu},$ (54)
which is motivated by an argument on the non-locality of entropy. We show that the gravity condensate with (46) saturates this exactly.
We first construct the spherically-symmetric static version of (54). From (48), we have $\tilde{s}_{\mu}dx^{\mu}=\tilde{s}_{t}dt$ with $\tilde{s}_{t}=-\frac{\sqrt{-g_{tt}}}{4\pi r^{2}}s$. We can calculate in (7)
$k^{\mu}k^{\nu}\nabla_{\mu}\tilde{s}_{\nu}$
$=k^{t}k^{r}(\nabla_{t}\tilde{s}_{r}+\nabla_{r}\tilde{s}_{t})$
$=k^{t}k^{r}\sqrt{-g_{tt}}\left(\frac{s}{4\pi r^{2}}\partial_{r}\log\sqrt{-g_{tt}}-\partial_{r}\left(\frac{s}{4\pi r^{2}}\right)\right).$ (55)
From this and (VI.2), (54) becomes
$\langle-T^{t}{}_{t}\rangle+\langle T^{r}{}_{r}\rangle$
$\geq\frac{\hbar}{2\pi}\frac{\sqrt{-g_{tt}}}{g_{rr}}\frac{|k^{t}|}{|k^{r}|}\left|\frac{s}{4\pi r^{2}}\partial_{r}\log\sqrt{-g_{tt}}-\partial_{r}\left(\frac{s}{4\pi r^{2}}\right)\right|$
$=\frac{\hbar}{2\pi}\left|\frac{\partial_{r}\log\sqrt{-g_{tt}}}{\sqrt{g_{rr}}}\frac{s}{4\pi r^{2}}-\frac{1}{\sqrt{g_{rr}}}\partial_{r}\left(\frac{s}{4\pi r^{2}}\right)\right|$ (56)
where we have applied (49). We then use the acceleration $\alpha=\frac{\partial_{r}\log\sqrt{-g_{tt}}}{\sqrt{g_{rr}}}$ to obtain the spherically-symmetric static version of (54):
$\left|\frac{\hbar\alpha}{2\pi}\frac{s}{4\pi r^{2}}-\frac{\hbar}{2\pi}\frac{1}{\sqrt{g_{rr}}}\partial_{r}\left(\frac{s}{4\pi r^{2}}\right)\right|\leq\langle-T^{t}{}_{t}\rangle+\langle T^{r}{}_{r}\rangle.$ (57)
Note that the entropy bound in this special case involves only the combination $\langle-T^{t}{}_{t}\rangle+\langle T^{r}{}_{r}\rangle$, independently of $\langle T^{\theta}{}_{\theta}\rangle$, and our evaluation of the entropy density in Sec.V.1.2 is consistent.
The interior metric (46) saturates (57) exactly at the leading order for $r\gg l_{p}$: The left hand side becomes $\frac{\hbar\alpha_{*}}{2\pi}\frac{s_{*}}{4\pi r^{2}}-\frac{\hbar}{2\pi}\frac{1}{\sqrt{g^{*}_{rr}}}\partial_{r}\left(\frac{s_{*}}{4\pi r^{2}}\right)\approx\frac{\hbar\alpha_{*}}{2\pi}\frac{s_{*}}{4\pi r^{2}}=\frac{1}{4\pi G\eta r^{2}}$ for $r\gg l_{p}$ (from (39) and (41)), while the right one becomes $\langle-T^{t}{}_{t}\rangle_{*}+\langle T^{r}{}_{r}\rangle_{*}=\frac{1}{4\pi G\eta r^{2}}$ (from (IV.3)). Thus, the gravity condensate saturates not only the global bound (5) but also the local condition (54) relevant to the local structure. This is so non-trivial that there must be something behind it.
VII 10 future prospects
The identity of a black hole is still unknown, and the microscopic constituents should involve fundamental degrees of freedom in quantum gravity, where there is no notion of classical spacetime geometry. Motivated by thermodynamics and (local) holography, in the present paper, we adopted the characterization that a black hole maximizes thermodynamic entropy for a given surface area. For spherical static configurations, we uniquely obtained the entropy-maximizing configuration consistent with local thermodynamics and the 4D semi-classical Einstein equation with many matter fields. That is, a self-gravitating collection of near-Planckian excited quanta condensates into the radially-uniform dense configuration without horizons or singularities as in Fig.2, where the self-gravity and large quantum pressure are balanced (called semi-classical gravity condensate). The interior metric (46) (say, for $n_{s}$ scalar fields, with the parameters (33)) satisfies the semi-classical Einstein equation self-consistently and non-perturbatively in $\hbar$. The bulk dynamics makes the entropy of the interior quanta exactly follow the Bekenstein-Hawking formula, leading to the Bousso bound for thermodynamic entropy and the saturation of the local sufficient conditions. This is a candidate picture of black holes in quantum theory.
Let us now discuss 10 future prospects and speculate on what possibilities there are for a more complete quantum description of black holes and a construction of quantum gravity.
1. Role of self-gravity in holography—. As shown in Sec.V.1.2, the self-gravity, represented by the self-consistent metric (32), changes the entropy (2) from the volume law to the area law Y1 . Then, what happens to quantum fields in the metric? From (41), the entropy per unit proper volume is given by $\frac{s_{*}(r)}{4\pi r^{2}}=\mathcal{O}\left(\frac{\sqrt{n}}{l_{p}r^{2}}\right)$. This is much smaller than the naive estimate without the effect of self-gravity, $\mathcal{O}\left(\frac{n}{l_{p}^{3}}\right)$, where all modes are assumed to be excited Bousso2 . This gap should be related to the self-consistent state $|\psi\rangle_{*}$ (described above (33) and (78)) in which only s-waves are excited and the other modes not. Indeed, the number of possible patterns of excited s-waves reproduces the Bekenstein-Hawking formula at a WKB-approximation level KY4 . Therefore, we can conjecture that the self-gravity suppresses the excitation of the local degrees freedom in the bulk and reduces the number of active bulk degrees of freedom, leading to the holographic property of entropy. This should be related to the remarkable fact that the local sufficient condition (54) for the Bousso bound is exactly saturated by the metric (32), which could be an essence of the bulk dynamics consistent with the holographic principle. For that, one could study quantum fields in the metric (32) by a semi-classical/perturbative Wheeler-De Witt equation Kiefer ; Kuchar ; Suvrat , since the Hamiltonian constraint $\mathcal{H}=0$ plays the key role in obtaining $S[g_{\mu\nu};R)$ in Sec.II.3.
2. Relation to other gravity-condensate models—. We discuss the relation to other gravity-condensate models and explore a possible path to a full-quantum formulation of black holes.
A model is in the framework of group field theory, a second quantization of loop quantum gravity Oriti1 . Gluing spherically-symmetric quantum-gravity states kinematically and maximizing the entropy for a given surface area, a quantum gravitational configuration appears with the entropy proportional to the surface area, and the coefficient is fixed by using the Unruh effect and thermodynamic relations. A remarkable point is that the holographic property of the entropy of the quanta living in the interior bulk holds for any size $r$, which corresponds to the radial uniformity in our semi-classical gravity condensate. (See Sec.8 of Ref.Y1 for more discussions.)
Another one is a view of black holes as Bose-Einstein condensates of gravitons Dvali3 . Introducing the occupation number $\mathcal{N}$ of gravitons in a gravitational field of a source with mass $M_{0}=\frac{a_{0}}{2G}$ and size $R$ by $\mathcal{N}=\frac{M_{0}a_{0}}{\hbar}\sim\frac{a_{0}^{2}}{l_{p}^{2}}(\leq\frac{R^{2}}{l_{p}^{2}})$ and maximizing it for a given size $R$, the black hole is characterized by $\mathcal{N}_{max}\sim\frac{R^{2}}{l_{p}^{2}}$.
To examine the relation to our gravity condensate, we first review how to obtain $\mathcal{N}\sim\frac{M_{0}a_{0}}{\hbar}$ Dvali3 . In the approximation of linear gravity, the gravitational energy of a source with size $R$ and mass $M_{0}$ can be estimated by $E_{g}\sim\frac{GM^{2}_{0}}{R}$, while the characteristic energy of a single graviton is given by $\epsilon_{g}\sim\frac{\hbar}{R}$. Therefore, we obtain the occupation number of gravitons as $\mathcal{N}\sim\frac{E_{g}}{\epsilon_{g}}\sim\frac{M_{0}a_{0}}{\hbar}$.
In the picture of Fig.3, we consider a spherical subsystem with a width $\Delta\hat{r}$ and the local energy $\Delta E_{loc}(r)=4\pi r^{2}\langle T^{\hat{t}\hat{t}}(r)\rangle\Delta\hat{r}$. For the condition (8), the approximation of linear gravity is valid within the subsystem. Following the above idea, then the gravitational energy can be estimated as $\Delta E_{g}(r)\sim\frac{G\Delta E_{loc}(r)^{2}}{\Delta\hat{r}}$, and the characteristic energy of a single graviton in the subsystem is given by $\epsilon_{g}(r)\sim\frac{\hbar}{\Delta\hat{r}}$. Therefore, the occupation number of gravitons within the subsystem is estimated by
$N_{g}(r)\sim\frac{\Delta E_{g}(r)}{\epsilon_{g}(r)}\sim\frac{G\Delta E_{loc}(r)^{2}}{\hbar}.$ (58)
In particular, for the case of our gravity condensate, we have $\Delta E_{loc}^{*}(r)\sim\sqrt{n}m_{p}$ (see around (28)) and get
$N_{g}^{*}(r)\sim n.$ (59)
Then, interesting similarities can be observed between $n$ and $\mathcal{N}$. For our condensate, the temperature $T_{loc}(r)\sim\frac{m_{p}}{\sqrt{n}}$, characteristic wavelength $\lambda_{*}(r)\sim\sqrt{n}l_{p}$, and energy $\Delta E_{loc}^{*}(r)\sim\sqrt{n}m_{p}$ are controlled by $n$. These manifestations of $n$ are exactly the same as those of $\mathcal{N}$ in the condensate of Ref.Dvali3 , although the former is for local subsystems and the latter for the entire system. Also, from (59), the condition $n\gg 1$ corresponds to that for the classicality in Ref.Dvali3 . Therefore, a possibility is that realizing the condensate in Ref.Dvali3 in local subsystems and connecting them consistently with the semi-classical Einstein equation lead to our gravity condensate.
Note that we have $N(r)\neq N_{g}(r)$ for a generic configuration, but our gravity condensate satisfies
$N^{*}(r)\sim n\sim N_{g}^{*}(r)$ (60)
from (27) and (59). This implies that matter quantum and gravity quantum in our gravity condensate play the same role in a sense, or they are indistinguishable as a quantum of some degree of freedom in quantum gravity. Then, the semi-classical gravity condensate would be a mixture of matter and gravity quanta, providing a basis of exploration of new degrees of freedom.
The two other models above discuss quantum gravitational effects without considering quantum matter ones, while our model describes matter quanta in the self-consistent (and non-perturbative) classical gravitational field. Therefore, a more detailed study of the connection between the three models would provide some clues to the above expectation.
3. Gravity-condensate phase—. We note that the gravity condensate, despite being self-gravitational, is uniform in the radial direction: (the leading values of) the local temperature $T_{loc}(r)$, the entropy density $s(r)$, and the 1D energy density $\rho_{1d}(r)$ are constant (see Sec.V.1.2). This implies Landau_SM that the gravity condensate is a kind of thermodynamic phase. (Let us call it the gravity-condensate phase). Indeed, this view works well in determining the position of the surface (45).
The phase could be quantum gravitational.
In general, for materials without self-gravity, quantum effects dominate the determination of macroscopic properties when the thermal wavelength $\lambda_{T}=\frac{\hbar}{\sqrt{mT}}$, a quantum length scale, is of the same order as the mean inter-particle distance $\rho_{N}^{-1/3}$, a classical length scale Landau_SM2 . Here, $m,~{}T,~{}\rho_{N}$ are mass of particles, temperature, and number density, respectively.
On the other hand, the relation (from (8))
$\lambda(r)\sim\frac{\hbar}{\epsilon(r)}\lesssim\mathcal{R}(r)^{-\frac{1}{2}}$ (61)
could be regarded as a condition under which the semi-classical approximation (1) holds: only when quanta have wavelengths $\lambda(r)$ shorter than the radius of spacetime curvatures $\mathcal{R}(r)^{-\frac{1}{2}}$ in the self-consistent gravitational field should the locality of quanta and spacetime be well-defined (at the resolution of $\sim\lambda(r)$) and the concept of the classical and continuum spacetime be established (as in Fig.3).
Here, noting the typical relation (10), we have a relativistic thermal wavelength $\lambda(r)\sim\frac{\hbar}{T_{loc}(r)}$. Therefore, we could expect the correspondence:
$\lambda_{T}\sim\rho_{N}^{-\frac{1}{3}}\overset{?}{\longleftrightarrow}\lambda(r)\sim\mathcal{R}(r)^{-\frac{1}{2}}.$ (62)
For the gravity condensate, (61) is saturated: $\lambda_{*}(r)\sim\mathcal{R}_{*}(r)^{-1/2}$ (from (35)), and also the equivalence in the number of gravity and matter quanta (60) holds. Thus, the correspondence (62) would imply that, the gravity condensate phase is governed essentially by quantum-gravitational effects while due to the large occupation number (60) for $n\gg 1$, the mean-field approximation of quantum gravity by (1) Kiefer holds albeit barely, leading to the semi-classical description by the metric (46).
A way to study this speculation is to first find an appropriate order parameter, study a Ginzburg-Landau-like theory or Gross–Pitaevskii-like equation including the effect of the redshift, and reproduce the radial uniformity or the interior metric. One could then consider the quantum many-body model behind it.
We finally discuss the notion of equilibrium in a self-gravitating system. In the absence of self-gravity, the most typical configuration maximizes the entropy for given macroscopic parameters and corresponds to the equilibrium state. Therefore, the derivation of the gravity condensate as the entropy-maximizing configuration suggests that the gravity condensate is the self-gravitating equilibrium state for a given surface area. This is consistent with the result of Appendix D.
Here, the effect of interactions should be important. In general, a phase is a uniform equilibrium state which is achieved through internal interactions Landau_SM . The parameter $\eta$ was originally introduced as a phenomenological parameter to represent the scattering/interaction effects of Hawking-like radiation by the gravitational potential and matters inside; a larger $\eta$ represents a larger interaction, whose macroscopic effect appears in the radial pressure $\langle T^{r}{}_{r}(r)\rangle$ in (IV.3) KY1 . Therefore, the above quantum many-body model should have interactions between matter and gravity quanta, which should explain $\eta$ and the large redshift $g_{tt}(r)\sim-e^{-\frac{R^{2}-r^{2}}{2\eta\sigma}}$ (or the time scale $\Delta t\sim\mathcal{O}(e^{a_{0}^{2}/l_{p}^{2}})$ in Sec.V.1) microscopically.
4. Path-integral evaluation of $S[g_{\mu\nu};R)$—. We have assumed the phenomenological form of thermodynamic entropy (2) for highly excited states and utilized local typicality and the Hamiltonian constraint to estimate the entropy $S[g_{\mu\nu};R)$, (16). This is a rough estimate but is the first one (in our knowledge) that includes non-perturbatively the effect of self-gravity consistent with the semi-classical Einstein equation for various configurations. However, this should be justified in a more microscopic manner.
A way to evaluate $S[g_{\mu\nu};R)$ field-theoretically is a path-integral method BY1 ; BY2 . For massless scalar fields in a self-consistent configuration $(g_{\mu\nu},|\psi\rangle)$, we should be able to use the propagator of massless particles restricted to a given size $R$ in the metric $g_{\mu\nu}$ and evaluate the density of states.
In the context of quantum gravity, a similar problem has been studied Jacobson_vol ; the dimension of the Hilbert space of a spatial region with a fixed proper volume is evaluated in the leading order saddle point approximation, where only gravity contribution is considered, to obtain the entropy-area law associated with the surface area of the saddle ball (see Bianca for a discrete model). On the other hand, motivated by (local) holography, we have fixed a surface area and considered typical configurations satisfying the semi-classical Einstein equation with matter fields, to find the entropy-maximizing one with the metric (46), leading to the Bekenstein-Hawking formula. Therefore, to understand the relation of the two clearly, one would consider a path integral in both gravity and matter for a fixed surface area and examine the relation between the saddle points and the self-consistent solutions of the semi-classical Einstein equation.
5. Thermodynamic entropy vs entanglement entropy—. Another field-theoretic approach to $S[g_{\mu\nu};R)$ would be to formulate it as entanglement entropy. In this case, the Unruh temperature is often applied to local Rindler regions as the local temperature Minic ; Casini ; Jacobson_entangle , but it does not necessarily agree with the thermodynamic temperature consistent with the self-consistent gravity. In the case of, say, self-gravitating thermal radiation, the Unruh temperature is different from the local temperature obtained by applying Stefan-Boltzmann law or Tolman’s law in the metric (see Appendix A). This is natural because the local temperature of an object in global thermodynamic equilibrium, which is fixed by Tolman’s law in the self-consistent metric, does not generically coincide with the Unruh temperature determined by the acceleration required to stay against the self-gravity. Therefore, the entanglement entropy calculated by such methods can be different from the thermodynamic entropy $S[g_{\mu\nu};R)$. The crucial point is whether to take into account the self-gravity determined by the semi-classical Einstein equation.
As argued in Sec.V.1.1, however, the local temperature (39) of the gravity condensate can be obtained both kinematically (from the Unruh formula) and dynamically (due to the particle creation). This is based on the self-consistent metric (46), where the relation $\lambda_{*}(r)\sim\mathcal{R}_{*}^{-\frac{1}{2}}(r)$ holds (possibly involving quantum gravity effects as mentioned in Prospect 3). Thus, we expect that applying such field-theoretic techniques locally enables us to evaluate the entropy density of the gravity condensate as entanglement entropy, and the entanglement entropy and the thermodynamic (Boltzmann) entropy agree and give the Bekenstein-Hawking formula. Indeed, this holds for the quantum gravity condensate proposed in Ref.Oriti1 . We would like to check this expectation in the future.
Note here that it is not clear whether entanglement entropy explains the Bekenstein-Hawking formula in general. As noted in footnote 1, the latter is a nontrivial function of gravitational charges that satisfies the first law of thermodynamics, and should be a type of thermodynamic entropy. Entanglement entropy, on the other hand, depends on the quantum state, and its leading value is fixed only by the surface area of the boundary and is not relevant to gravitational charges directly.
Furthermore, the origin of the Bekenstein-Hawking entropy should be consistent with a physical understanding of information problem (see below): through the unitary evolution, the microscopic states contributing to the entropy should correspond one-to-one to the initial states of a collapsing matter forming the black hole.
6. Physical resolution of the information problem—. The purification of the initial Minkowski vacuum after evaporation and the “formal” derivation of the Page curve are only part of the information problem. More essentially, we need to understand the dynamical mechanism by which the initial wavefunction of a collapsing matter forming the black hole is recovered after evaporation, leading to the physical Page curve as the result of the unitary evolution of matter and gravity. We must also clarify the interior structure with quantum dynamics that resolves the singularity. Such overall consistency will show the true identity of the black hole.
In the semi-classical gravity condensate, the information of a (typical) collapsing matter is stored in the bulk interior, and the amount agrees with the Bekenstein-Hawking formula. The condensate has neither horizon nor singularity (although we still need to understand the small center part at a fully quantum-gravity level), and almost all parts evaporate in the vacuum due to the Hawking-like radiation. Therefore, it should be natural to expect that (most of) the information recovers after the evaporation. As mentioned above, however, we still need to clarify the mechanism consistent with the energy flow: how the information of the matter leaks out gradually during the evaporation. A possibility is that scattering between a collapsing matter and the radiation occurs frequently at each point inside KY2 , and such interactions should transfer the information to the emitted radiation, reproducing the physical Page curve.
To see this explicitly, one would at least need to check the initial-state dependence of the radiation emitted after the interaction. However, the condensate is in a typical state, and such dependence is not easy to see. One idea to overcome this is to consider perturbations from the typical state, adding a small atypical portion to the condensate and tracking it during evaporation. Another is to study how to distinguish between two typical states in a certain protocol and implement it in this model.
7. Non-typical black holes—. The semi-classical gravity condensate has been obtained as the entropy-maximizing configuration for a given surface area and thus should be the most typical black hole. As described around Fig.4, it can be formed by adiabatic processes in a heat bath of Hawking temperature. Then, what is a non-typical black hole that is formed by a generic collapse of matter not by such a reversible process?
A way to consider it is to analyze the time evolution of a collapsing matter including the backreaction from particle creation during the collapse (see footnote 15). Indeed, we can consider the matter as consisting of many spherical shells with small energy, distributed depending on the initial configuration and solve (1) self-consistently KY2 . It shows that the dense structure with the metric (46) is formed only around the surface, while the structure in deeper regions depends on the details of the initial distribution. The entropy is smaller than (42), and this is a non-typical black hole.
Here, it is important to consider a finite width even when discussing collapse of a spherical shell. For simplicity it is often modeled by an infinitely thin shell BD ; BHmodel , but physically, any collapsing matter is an excitation of quantum fields with a physical information $|\psi\rangle$, which cannot be localized completely in the radial direction. Considering a finite-width shell as a collection of many tiny shells and including the evaporation effect, we can see that the tangential pressure $\langle T^{\theta}{}_{\theta}\rangle$ occurs dynamically and the gravity-condensate structure appears only around $r=a_{0}$ while the interior is vacuum KY2 ; KY4 . This is a result of the 4D dynamics; in the 4D conservation law $\nabla_{\mu}\langle T^{\mu}{}_{\nu}\rangle=0$, which contains $\langle T^{\theta}{}_{\theta}\rangle$, it can be large due to the 4D Weyl anomaly KY3 , while in the 2D cases, there is no tangential direction.
Another way is to consider a solution space, consistent with the 4D Weyl anomaly, of the semi-classical Einstein equation and study various configurations in a non-perturbative manner for $\hbar$ HKLY . It shows that the most typical ones for a given surface area are similar to the semi-classical gravity condensate, and the number of such configurations explains the Bekenstein-Hawking entropy (except for $O(1)$ numerical factor). Also, non-typical ones have the dense structure only around the surface, which is consistent with the one above.
Therefore, these different studies based on non-perturbative dynamics of the semi-classical Einstein equation give almost the same picture of the typical and non-typical black holes. It is interesting to study a time-dependent perturbation and consider how such a non-typical one evolves to the typical one.
8. Relation to the classical picture of black holes—. One might wonder why the gravity condensate is so different from the classical picture of black holes, a vacuum region surrounded by a horizon. First, the classical one was originally derived from the classical dynamics of a collapsing matter and has not been confirmed observationally yet. Second, even at the semi-classical level, infinitely thin-shell models with 2D approximation are often used, leading to almost the same picture as the classical one BHmodel . In general, however, quantum fluctuations of modes with arbitrary angular momentum are induced in 4D spherically symmetric spacetime (which leads to 4D Weyl anomaly, for example). As mentioned above, the consideration of a finite width is important in the information problem. Furthermore, the backreaction from evaporation during the collapse is often neglected, but must be included (see footnote 15). On the other hand, our argument takes these points into account and obtains the picture of the gravity condensate. Thus, the difference appears.
Another aspect in the difference between the conventional picture and the gravity condensate is that they belong to different branches in a solution space of the semi-classical Einstein equation. In Ref.HKLY , we constructed a self-consistent equation of the energy distribution $\frac{a(r)}{2G}$ including the effect of the 4D Weyl anomaly, and examined the structure of the solution space in a non-perturbative manner for $\hbar$. We then found that there exist two branches: one is perturbative and contains Schwarzschild-like metrics, while the other is non-perturbative and includes the dense solution (32). The point is that, the higher derivative term $\Box R_{RicciScalar}$, which appears generically in the anomaly BD , causes transitions between the two branches as $r$ changes. Therefore, it would be interesting to study 4D semi-classical time evolution of a collapsing matter including such higher derivative effects and see how the gravity condensate is generically formed.
9. Gravitational field with finite entropy—. We briefly comment on entropy in gravity. An interesting aspect in estimating the entropy $S[g_{\mu\nu};R)$ in Sec.II is that local typicality and Hamiltonian constraint assign a finite entropy to a gravitational field $g_{\mu\nu}$. This is reminiscent of two things. First, it is similar to the idea that the Einstein equation corresponds to the equation of state in spacetime thermodynamics Jacobson ; Pad . Second, it should be relevant to the view of entropy as a gravitational charge Wald ; SY (see Sec.8 of Ref.Y1 for details). These could be relevant each other.
10. Phenomenology—. We finally discuss the phenomenological aspect of the semi-classical gravity condensate. In Ref.CY , an investigation of imaging of the gravity condensate showed that, despite the absence of an event horizon, the image is significantly darkened by the strong redshift of (3) and almost identical to the classical black-hole image, giving the consistency with the current data. Furthermore, the intensity around the inner shadow is slightly enhanced when the emission is a bit inside the surface, which may be a future observable prediction for characterizing the condensate. It is also interesting to investigate gravitational waves in this model; in particular, some echo signal can be expected due to the existence of the surface structure echo . For a more realistic phenomenology, furthermore, it should be important to generalize the gravity condensate to a rotating case (see Ref.KY2 for a slowly-rotating case).
Acknowledgments
Y.Y. thanks C.Barcelo, F.Becattini, R.Casadio, C.Y.Chen, C.Goeller, T.Harada, C.Kelly, E.Livine, N.Nakagawa, A.Pesci, and Y.Sakatani for inspiring discussions and valuable comments. Y.Y. is partially supported by Japan Society of Promotion of Science (Grants No.21K13929) and by RIKEN iTHEMS Program.
Appendix A Self-gravitating thermal radiation
To demonstrate the self-gravity dependence of entropy explicitly in an example, we provide a review for the entropy of self-gravitating thermal radiation in a different manner from Ref.Sorkin . We also check its consistency to $S[g_{\mu\nu};R)$ and the upper bound.
A.1 Metric
We consider a spherically-symmetric static configuration of self-gravitating ultra-relativistic fluid with size $R$, mass $M_{0}=\frac{a_{0}}{2G}$, and equation of state $\rho=3p$, where $\rho(r)\equiv\langle-T^{t}{}_{t}(r)\rangle$ and $p(r)\equiv\langle T^{r}{}_{r}(r)\rangle=\langle T^{\theta}{}_{\theta}(r)\rangle$. For simplicity, we here neglect a small contribution from Weyl anomaly from the curvatures BD and interactions fluidbook . We construct its interior metric, for (7), in a heuristic manner. (See Weinberg for another derivation.)
First, the equation of state $\rho=3p$, equivalent to $\langle T^{\mu}{}_{\mu}\rangle=0$, means that there is no special length scale in the system. Therefore, the order of the magnitude of the curvature at $r$ should be $\mathcal{R}(r)\sim\frac{1}{r^{2}}$. Also, $\rho=3p$ indicates $\rho\sim p$, and both $\rho$ and $p$ contribute to the curvature almost equally. Thus, from the Einstein equation (1), we have $\rho(r)\sim\frac{1}{Gr^{2}}$, which can be expressed as $\rho(r)=\frac{1-C}{8\pi Gr^{2}}$ with a $\mathcal{O}(1)$ constant $C$. Applying (13) to this, we have
$\frac{a(r)}{2G}=\frac{1-C}{2G}r.$ (63)
This must be positive due to radiation excitation and must be smaller than $\frac{4r}{9G}$ from Buchdahl’s limit Buchdahl ; Weinberg , which requires $\frac{1}{9}<C<1$. Then, the interior metric (7) becomes
$ds^{2}=-Ce^{A(r)}dt^{2}+\frac{1}{C}dr^{2}+r^{2}d\Omega^{2}.$ (64)
Next, $A(r)$ is determined from the ultra-relativistic fluid condition: $\langle-T^{t}{}_{t}\rangle=3\langle T^{r}{}_{r}\rangle$ and $\langle T^{r}{}_{r}\rangle=\langle T^{\theta}{}_{\theta}\rangle$. (64) gives
$-G^{t}{}_{t}$ $=\frac{1-C}{r^{2}},$ (65)
$G^{r}{}_{r}$ $=-\frac{1-C}{r^{2}}+\frac{CA^{\prime}(r)}{r},$ (66)
$G^{\theta}{}_{\theta}$ $=\frac{C}{2r}A^{\prime}(r)+\frac{1}{4}CA^{\prime}(r)^{2}+\frac{1}{2}CA^{\prime\prime}(r).$ (67)
Using these and (1), $\langle-T^{t}{}_{t}(r)\rangle=3\langle T^{r}{}_{r}(r)\rangle$ leads to $A(r)=A_{0}+\frac{4(1-C)}{3C}\log r$ ($A_{0}$: constant). Applying this $A(r)$, (66) and (67) to $\langle T^{r}{}_{r}(r)\rangle=\langle T^{\theta}{}_{\theta}(r)\rangle$ through (1), we have $(1-C)(-4+7C)=0$, giving $C=\frac{4}{7}$, which is consistent with Buchdahl’s limit. Then, we get $A(r)=A_{0}+\log r$, and (64) becomes
$ds^{2}=-\frac{4}{7}e^{A_{0}}rdt^{2}+\frac{7}{4}dr^{2}+r^{2}d\Omega^{2},$ (68)
and the energy density is given by Weinberg
$\rho(r)=\frac{3}{56\pi Gr^{2}}.$ (69)
Here, the size $R$ is determined by applying (63) with $C=\frac{4}{7}$ and $a(R)=a_{0}$:
$R=\frac{7}{3}a_{0}.$ (70)
Finally, we connect (68) to the Schwarzschild metric with mass $\frac{a_{0}}{2G}$. This requires the continuity of $-g_{tt}(r)$ at $r=R$ Poisson : $-g_{tt}(R)=\frac{4}{7}e^{A_{0}}R=1-\frac{a_{0}}{R}$, leading through (70) to $e^{A_{0}}=\frac{1}{R}$. Thus, we reach the final form:
$ds^{2}=-\frac{4}{7}\frac{r}{R}dt^{2}+\frac{7}{4}dr^{2}+r^{2}d\Omega^{2}~{}~{}~{}{\rm for}~{}~{}l_{p}\ll r\leq R.$ (71)
Note that the curvatures are small for $r\gg l_{p}$: $R_{\mu\nu\alpha\beta}R^{\mu\nu\alpha\beta}\sim\frac{1}{r^{4}}$. Therefore, the contribution from 4D Weyl anomaly BD is indeed small for $r\gg l_{p}$, where our approximation is valid.
A.2 Entropy
Now, let us evaluate the entropy. First, the equation of state $\rho=3p$ and thermodynamics lead to the Stefan-Boltzmann law:
$\rho(r)=\frac{kn_{f}}{4\hbar^{3}}T_{loc}(r)^{4},$ (72)
where $k$ is a $\mathcal{O}(1)$ numerical constant, and $n_{f}$ is the number of the degrees of freedom of radiation Landau_SM . This and (69) determine the local temperature at $r$:
$T_{loc}(r)\sim\frac{n_{f}^{-\frac{1}{4}}\hbar}{\sqrt{l_{p}r}}.$ (73)
Note that this agrees with one obtained by applying Tolman’s law Landau_SM to the metric (71), $\frac{T_{0}}{\sqrt{-g_{tt}(r)}}\sim r^{-\frac{1}{2}}$, and thus the metric (71) is consistent with thermodynamics and general relativity. From (10), this gives the characteristic excitation $\epsilon(r)\sim T_{loc}(r)$ for a typical state $|\psi\rangle$.
Using $\rho=3p$, the Gibbs relation $\rho+p=T_{loc}s_{3d}$ Groot , (72) and (73), we obtain the entropy per unit proper volume, $s_{3d}(r)\equiv\frac{s(r)}{4\pi r^{2}}$, as
$s_{3d}(r)=\frac{kn_{f}}{3\hbar^{3}}T_{loc}(r)^{3}\sim\frac{n_{f}^{\frac{1}{4}}}{(l_{p}r)^{\frac{3}{2}}}.$ (74)
Thus, applying (74) and (71) to (2), we get Sorkin
[…]
which is different from the usual volume law $\sim R^{3}$. This is a consequence of the self-gravity dependence of entropy.
A.3 Consistency check
We use this example to examine the consistency of our argument on $S[g_{\mu\nu};R)$ and the upper bound.
First, the ratio between (72) and (74) leads to
$s_{3d}(r)=\frac{4\langle-T^{t}{}_{t}(r)\rangle}{3T_{loc}(r)},$ (76)
which gives a consistency to the estimation (11).
Second, we check the conditions (8) and (12). We can use $\lambda(r)\sim\frac{\hbar}{\epsilon(r)}$, (10) and (73) to get $\lambda(r)\sim n_{f}^{\frac{1}{4}}\sqrt{l_{p}r}$, which is shorter than $\mathcal{R}(r)^{-\frac{1}{2}}\sim r$ (from (69)) for $r\gg l_{p}$, and the condition (8) holds. Then, from (8) and (74), we can estimate (12) as
$n_{f}^{\frac{1}{2}}\frac{r}{l_{p}}\lesssim N(r)\lesssim n_{f}^{\frac{1}{4}}\frac{r^{\frac{3}{2}}}{l_{p}^{\frac{3}{2}}},$ (77)
which is large for $r\gg l_{p}$.
Third, we study the relation to the second law. Locally $\frac{\partial\tilde{s}(r)}{\partial T_{loc}(r)}>0$ holds from (74), while for the whole part, we have $\frac{\partial S}{\partial T_{0}}<0$ indicating that the total heat capacity is negative Landau_SM , where we use (75) and $T_{0}=T_{loc}(R)\propto R^{-1/2}$ (from (73)). Thus, due to the long-range nature of gravity, the thermodynamic behavior is different depending on the region being considered.
Fourth, we check that thermal radiation does not saturate the sufficient conditions for the Bousso bound. We first study (19). The right hand side is evaluated as $\frac{1}{\hbar}\sqrt{g_{rr}(r)}(r-a(r))4\pi r^{2}\langle-T^{t}{}_{t}(r)\rangle\sim\frac{r}{l_{p}^{2}}$ from (69) and (71), while the left hand side as $s=4\pi r^{2}s_{3d}\sim n_{f}^{\frac{1}{4}}\frac{r^{\frac{1}{2}}}{l_{p}^{\frac{3}{2}}}$ from (74). Therefore, (19) is not saturated for $r\gg l_{p}$ and not so large $n_{f}$.
We next examine (57). The right hand side becomes $\rho+p=\frac{4}{3}\rho=\frac{1}{14\pi Gr^{2}}$ from (69). For the left hand side, we have $\alpha(r)=\frac{\partial_{r}\log\sqrt{-g_{tt}(r)}}{\sqrt{g_{rr}(r)}}\sim\frac{1}{r}$ from (71) and obtain from (74) $\left|\frac{\hbar\alpha}{2\pi}\frac{s}{4\pi r^{2}}-\frac{\hbar}{2\pi}\frac{1}{\sqrt{g_{rr}}}\partial_{r}\left(\frac{s}{4\pi r^{2}}\right)\right|\sim\frac{n_{f}^{\frac{1}{4}}}{Gr^{2}}\sqrt{\frac{l_{p}}{r}}$. Thus, (57) is not saturated.
Appendix B Self-consistency of $(g_{\mu\nu}^{*},|\psi\rangle_{*})$
We give a short review for the derivation of the self-consistent values (33) of $(\sigma,\eta)$ where the origin of the tangential pressure can be seen (see Ref.KY4 for details). This is also a demonstration of our self-consistent argument in the semi-classical Einstein equation (1).
We start with a review about how to solve the semi-classical Einstein equation (1) self-consistently. First, we consider the physical system and problem of interest and construct a candidate metric $g_{\mu\nu}$ for it, say, by making a model or a thought experiment. Next, using that metric as a background spacetime, we examine the behavior of the matter fields and identify a candidate state $|\psi\rangle$ for the system. Then, we use the solutions of the matter field equations, construct the regularized energy-momentum tensor, and renormalize it to remove divergences. Finally, we equate the obtained renormalized energy-momentum tensor $\langle\psi|T_{\mu\nu}|\psi\rangle$ with the Einstein tensor $G_{\mu\nu}$ calculated from the metric and solve (1). If it can be solved consistently, the candidate $(g_{\mu\nu},|\psi\rangle)$ is the self-consistent solution. If not, we set up another candidate and repeat the procedure. Note that this self-consistent analysis of (1) allows us to obtain a non-perturbative solution in $\hbar$.
Let’s implement this program by setting (32) for the inner region (28) and the Schwarzschild metric with $a_{0}$ for the outer one $R\leq r$ as the candidate metric.
(1) Candidate state. It should be natural to consider its formation process and find a candidate state for the gravity condensate. As discussed in Sec.IV.3.1, we can form it slowly in a heat bath of Hawking temperature $\sim\frac{\hbar}{a_{0}}$. Studying each mode of scalar field, $\phi(x)=e^{-i\omega t}\varphi(r)Y_{lm}(\theta,\phi)$, in the formation, we can find that there exist two types of modes inside. One is bound modes with various angular momenta $l$, which are trapped inside the condensate and cannot be excited due to a constraint from the Bohr-Sommerfeld quantization condition. The other is a continuum mode of s-waves, which can go to and from the outside and can be excited with $\omega\sim\frac{1}{a_{0}}$ to express thermal radiation from the bath and the Hawking radiation produced inside. Therefore, we set a candidate state $|\psi\rangle_{*}$ such that
$\langle\psi|T_{\mu\nu}|\psi\rangle_{*}\approx\langle 0|T_{\mu\nu}|0\rangle+T_{\mu\nu}^{(\psi)}.$ (78)
The first term is the contribution from the vacuum fluctuation of all modes in the ground state $|0\rangle$ in (32), and the second one is that from the excitation of the continuum mode of s-waves. We here assume that the second contribution is so excited that it can be approximated as a classical one.
(2) Regularization. To regularize the energy-momentum operator, we use the dimensional regularization scheme. In a $d=4+\epsilon$ dimensional spacetime, the semi-classical Einstein equation coupled with $n_{s}$ free massless scalar fields can be expressed as a regularized operator equation BD :
$G_{\mu\nu}=8\pi G\left[\mu^{-\epsilon}T_{\mu\nu}-2\left(\frac{\hbar n_{s}}{1152\pi^{2}\epsilon}+\alpha(\mu)\right)H_{\mu\nu}\right.$
$\left.-2\left(-\frac{\hbar n_{s}}{2880\pi^{2}\epsilon}+\beta(\mu)\right)K_{\mu\nu}-2\left(\frac{\hbar n_{s}}{2880\pi^{2}\epsilon}+\gamma(\mu)\right)J_{\mu\nu}\right].$ (79)
Here, $\hbar\mu$ is a renormalization point, $T_{\mu\nu}$ is the regularized energy-momentum tensor operator, and the other tensors are proportional to the identity operator. The counter terms with $1/\epsilon$ is chosen as those required by the minimal subtraction scheme, and each tensor is defined as, respectively, $H_{\mu\nu}\equiv\frac{1}{\sqrt{-g}}\frac{\delta}{\delta g^{\mu\nu}}\int d^{d}x\sqrt{-g}R^{2}$, $K_{\mu\nu}\equiv\frac{1}{\sqrt{-g}}\frac{\delta}{\delta g^{\mu\nu}}\int d^{d}x\sqrt{-g}R_{\alpha\beta}R^{\alpha\beta}$, and $J_{\mu\nu}\equiv\frac{1}{\sqrt{-g}}\frac{\delta}{\delta g^{\mu\nu}}\int d^{n}x\sqrt{-g}R_{\alpha\beta\gamma\delta}R^{\alpha\beta\gamma\delta}$. We also have the renormalized coupling constants at energy scale $\hbar\mu$:
$\alpha(\mu)$ $=\alpha_{0}-\frac{\hbar n_{s}}{2304\pi^{2}}\log\left(\frac{\mu^{2}}{\mu_{0}^{2}}\right),~{}$
$\beta(\mu)$ $=\beta_{0}+\frac{\hbar n_{s}}{5760\pi^{2}}\log\left(\frac{\mu^{2}}{\mu_{0}^{2}}\right),~{}$
$\gamma(\mu)$ $=-\frac{\hbar n_{s}}{5760\pi^{2}}\log\left(\frac{\mu^{2}}{\mu_{0}^{2}}\right).$ (80)
Here, $\alpha_{0}$ and $\beta_{0}$ fix a 4D theory at energy scale $\hbar\mu_{0}$ while we have chosen $\gamma_{0}=0$ because of the 4D Gauss-Bonnet theorem. Thus, we obtain a more explicit form of the renormalized energy-momentum tensor in the right hand side of (1):
$\langle\psi|T_{\mu\nu}|\psi\rangle=\lim_{\epsilon\to 0}\left[\mu^{-\epsilon}\langle\psi|T_{\mu\nu}|\psi\rangle_{reg}\right.$
$-\frac{\hbar n_{s}}{1440\pi^{2}\epsilon}\left(\frac{5}{2}H_{\mu\nu}-K_{\mu\nu}+J_{\mu\nu}\right)$
$\left.-2(\alpha(\mu)H_{\mu\nu}+\beta(\mu)K_{\mu\nu}+\gamma(\mu)J_{\mu\nu})\right].$ (81)
(3) Matter fields. We now focus on $\langle 0|T_{\mu\nu}|0\rangle$ in (78), which will lead to the self-consistent value of $\sigma$. To evaluate $\langle 0|T_{\mu\nu}|0\rangle_{reg}$, we solve the matter field equation $\Box\phi=0$ in the $(4+\epsilon)$-dimensional spacetime manifold ×$\mathcal{M}\times\mathbb{R}^{\epsilon}$, where $\mathcal{M}$ is our 4D physical spacetime (32) and $\mathbb{R}^{\epsilon}$ is $\epsilon$-dimensional flat spacetime KN :
$ds^{2}=-\frac{2\sigma}{r^{2}}e^{\frac{r^{2}}{2\sigma\eta}}dt^{2}+\frac{r^{2}}{2\sigma}dr^{2}+r^{2}d\Omega^{2}+\sum_{a=1}^{\epsilon}(dy^{a})^{2}.$ (82)
Using the fact that (32) is locally $AdS_{2}\times S^{2}$ (with AdS radius $L\equiv\sqrt{2\eta^{2}\sigma}$ and $S^{2}$ radius $r$), we can solve $\Box\phi=0$ around a point $r=r_{0}$ perturbatively by a $1/r$ expansion for $r\gg L$. Then, the 0-th order solution for the bound modes is given by
$\phi(x)=\sum_{i}(a_{i}u_{i}^{(0)}(x)+a_{i}^{\dagger}u_{i}^{(0)\ast}(x)),$ (83)
where $a_{i}|0\rangle=0$, $\sum_{i}=\sum_{n,l,m}\int d^{\epsilon}k$, and
$u_{i}^{(0)}(t,r,\theta,\phi,y^{a})=$
$\sqrt{\frac{\hbar\eta}{2r_{0}}}\sqrt{\frac{\partial\omega_{nl}}{\partial n}}e^{-i\omega_{nl}t}e^{-\frac{r^{2}}{8\sigma\eta}}J_{j}(X)Y_{lm}(\theta,\phi)\frac{e^{ik\cdot y}}{(2\pi)^{\epsilon/2}}.$ (84)
Here, $n$ represents the quantum number satisfying the Bohr-Sommerfeld quantization condition; the factors in front of $e^{-i\omega_{nl}t}$ is the normalization consistent with the condition and commutation relation; and $J_{j}(X)$ is the Bessel function, where $j\equiv\sqrt{L^{2}(l(l+1)/r_{0}^{2}+k^{2})+\frac{1}{4}}$ and $X\equiv Le^{\frac{r_{0}(r_{0}-r)}{2\sigma\eta}}\omega/\sqrt{-g_{tt}(r_{0})}$.
(4) Renormalization. After a long calculation using (83), we can obtain the leading term of the renormalized energy-momentum tensor (B):
$\langle 0|T^{\mu}{}_{\nu}|0\rangle^{(0)}=$ $\left(\begin{array}[]{cccc}1&&&\\ &1&&\\ &&-1&\\ &&&-1\end{array}\right)\langle 0|T^{t}{}_{t}|0\rangle^{(0)}$ (89)
$+\left(\begin{array}[]{cccc}0&&&\\ &0&&\\ &&1&\\ &&&1\end{array}\right)\frac{\hbar n_{s}}{1920\pi^{2}\eta^{4}\sigma^{2}},$ (94)
where the components are in the order of $(t,r,\theta,\phi)$ and
$\langle 0|T^{t}{}_{t}|0\rangle^{(0)}$
$=\frac{\hbar n_{s}}{1920\pi^{2}\eta^{4}\sigma^{2}}\left[2c+\gamma+\log\frac{1}{32\pi\eta^{2}\sigma\mu^{2}_{0}}-\frac{960\pi^{2}}{\hbar n_{s}}(2\alpha_{0}+\beta_{0})\right].$ (95)
Here, $\gamma$ is Euler’s constant and $c$ is the non-trivial finite value for $|0\rangle$: $c=0.055868$. We note that the trace part
$\langle 0|T^{\mu}{}_{\mu}|0\rangle^{(0)}=\frac{\hbar n_{s}}{960\pi^{2}\eta^{4}\sigma^{2}}$ (96)
is independent of $(\alpha_{0},\beta_{0})$. Here, we can use (34) and check that this agrees with the leading value of the 4D Weyl anomaly BD ; Nicolai :
$\langle T^{\mu}{}_{\mu}\rangle_{Anomaly}=\frac{\hbar n_{s}}{2880\pi^{2}}\left(R_{\mu\nu\alpha\beta}^{2}-R_{\mu\nu}^{2}+\frac{5}{2}R_{RicciScalar}^{2}\right).$ (97)
(5) Self-consistent solution. Finally, we use (34), (78) and (96) and solve the trace part of (1) at the leading order:
$(G^{\mu}{}_{\mu})^{(0)}$ $=8\pi G\langle\psi|T^{\mu}{}_{\mu}|\psi\rangle^{(0)}$
$\Rightarrow\frac{1}{\sigma\eta^{2}}$ $=8\pi G\frac{\hbar n_{s}}{960\pi^{2}\eta^{4}\sigma^{2}}$
$\Rightarrow\sigma$ $=\frac{n_{s}l_{p}^{2}}{120\pi\eta^{2}},$ (98)
which gives the self-consistent value of $\sigma$ in (33). Here, we have dropped the contribution from $T^{(\psi)\mu}{}_{\mu}$ because the mode integrations over $\omega$ and $l$ lead to $\mathcal{O}(1)$ terms in (89), and $T^{(\psi)\mu}{}_{\mu}$, including only s-waves, cannot produce a $\mathcal{O}(1)$ term. Note that (B) is a non-perturbative argument in that it does not hold in the limit $\hbar\to 0$.
Furthermore, we can consider $T_{\mu\nu}^{(\psi)}=\mathcal{O}(r^{-2})$ and calculate the 1st order contribution $\langle 0|T^{\mu}{}_{\mu}|0\rangle^{(1)}=\mathcal{O}(r^{-2})$ to find the self-consistent value of $\eta$ in a similar manner. Here, we need to choose a theory with $(\alpha_{0},\beta_{0})$ such that $\langle 0|T^{t}{}_{t}|0\rangle^{(0)}=0$ in (B) and $21\leq\eta<$ 2 in (33) hold. We thus conclude that $(g_{\mu\nu}^{*},|\psi\rangle_{*})$ with (33) satisfies the semi-classical Einstein equation (1) self-consistently and non-perturbatively in $\hbar$.
Here, we can see explicitly the origin of the large tangential pressure $\langle\psi|T^{\theta}{}_{\theta}|\psi\rangle^{*}$ in (IV.3). Under the condition that $\langle 0|T^{t}{}_{t}|0\rangle^{(0)}=0$, the second term in (89) gives
$\langle 0|T^{\theta}{}_{\theta}|0\rangle^{(0)}=\frac{\hbar n_{s}}{1920\pi^{2}\eta^{4}\sigma^{2}}=\frac{15}{2Gn_{s}l_{p}^{2}},$ (99)
where we have used (B). This is the half of the Weyl anomaly (96), which is given by the curvatures (97). Therefore, the pressure originates from 4D quantum fluctuations induced by the curved spacetime (32).
Appendix C Maximum entropy from uniformity
We provide another derivation of (24): we derive it from a radially-uniform condition, instead of setting $\epsilon=\epsilon_{max}$. This indicates that the radial uniformity is a sufficient condition for entropy maximization; since the converse proposition of this is given in Sec.IV.1, we can therefore conclude that a necessary and sufficient condition for entropy maximization is radial uniformity. This would be natural for a spherically symmetric system in equilibrium according to thermodynamics in flat space, but it is non-trivial for a self-gravitating system.
To express radial uniformity, we use the occupation number (12) and impose
$N(r)={\rm const}.\equiv N_{0}~{}~{}{\rm for}~{}\Delta\hat{r}\sim\lambda(r),$ (100)
where $N_{0}=\mathcal{O}(1)\gg 1$. This means maximum uniformity, in the sense that the number of excited quanta is constant even in subsystems with the smallest width.
We first prepare a useful formula for analyzing (100). From the discussion below (12), we have $N(r)\sim\frac{\Delta E_{loc}}{\epsilon}\sim\frac{4\pi r^{2}\hbar\langle-T^{t}{}_{t}\rangle}{\epsilon^{2}}$ for $\Delta\hat{r}\sim\lambda\sim\frac{\hbar}{\epsilon}$. Applying (13), we obtain
$N(r)\sim\frac{m_{p}^{2}}{\epsilon(r)^{2}}\partial_{r}a(r).$ (101)
Setting (100), we then combine (101) and (23) to get
$N_{0}l_{p}^{2}\sim r(r-a_{*}(r))\partial_{r}a_{*}(r).$ (102)
Here, $N(r)={\rm const.}$ means through (101) that $\epsilon(r)>0$ and $\partial_{r}a_{*}(r)>0$. Hence, we can solve $a_{*}=a_{*}(r)$ for $r$ to have $r=r(a)$ (here we write $a=a_{*}$ for simplicity) and express (102) as
$\frac{dr(a)}{da}=\frac{1}{\gamma}r(a)(r(a)-a),$ (103)
where $\gamma={\rm const.}=\mathcal{O}(N_{0}l_{p}^{2})$.
The general solution to (103) is given by
$r(a)=\frac{e^{-x^{2}}}{C-\sqrt{\frac{\pi}{2\gamma}}{\rm erf}(x)},$ (104)
[…]
$r(a)\approx\frac{e^{-x^{2}}}{\sqrt{\frac{\pi}{2\gamma}}-C+e^{-x^{2}}\left[\frac{1}{\sqrt{2\gamma}x}-\frac{1}{2\sqrt{2\gamma}x^{3}}+\mathcal{O}(x^{-5})\right]}.$ (105)
If $\sqrt{\frac{\pi}{2\gamma}}-C\neq 0$, we would have $r(a)\sim e^{-x^{2}}$ for $x\gg 1$, but it is not consistent with $r,a\gg l_{p}$. Therefore, we must have $C=\sqrt{\frac{\pi}{2\gamma}}$. Then, we obtain
$r(a)$ $\approx\sqrt{2\gamma}x+\sqrt{\frac{\gamma}{2}}\frac{1}{x}+\mathcal{O}(x^{-3})$
$=a+\frac{\gamma}{a}+\mathcal{O}(a^{-3}),$ (106)
which means for $a\gg l_{p}$
$a_{*}(r)\approx r-\frac{\gamma}{r}.$ (107)
We now substitute (107) into (101) to get
$\epsilon(r)\sim\frac{m_{p}}{\sqrt{N_{0}}}.$ (108)
Comparing this to (17), we find $N_{0}\sim n$. Therefore, (107) means (24).
Appendix D Tolman’s law does not lead to maximum entropy.
We show that keeping Tolman’s law under the semi-classical condition (17) and consistency with local thermodynamics does not yield maximum entropy for a given surface area. This gives an explanation for the violation of Tolman’s law to obtain maximum entropy, discussed in Sec.V.1.1. Note that this is consistent with Refs.Green ; Xia , where isotropic fluid is assumed and quantities other than a surface area are fixed (see also footnote 9).
Suppose that, instead of setting $\epsilon(r)=\epsilon_{max}$ at each $r$, we use Tolman’s law $T_{loc}(r)\sqrt{-g_{tt}(r)}={\rm const}.$ consistent with the semi-classical condition (17) and saturate the entropy bound (21). From (10), we have
$\epsilon(r)=\frac{\sqrt{-g_{tt}(r_{0})}}{\sqrt{-g_{tt}(r)}}\epsilon_{max},$ (109)
where the maximum excitation $\epsilon_{max}\sim\frac{m_{p}}{\sqrt{n}}$ is reached at an innermost radius $r_{0}$. Then, (109) should saturate the local condition (18): $\frac{\hbar}{\epsilon(r)}\sim\sqrt{g_{rr}(r)}(r-a(r))=\frac{r}{\sqrt{g_{rr}(r)}}$, since it has led to (21). Taking the square of the both hands and applying (109), we get
$(-g_{tt}(r))g_{rr}(r)=Cr^{2},$ (110)
where $C$ is a positive constant. Therefore, the interior metric is given by
$ds^{2}=-\frac{Cr^{2}}{g_{rr}(r)}dt^{2}+g_{rr}(r)dr^{2}+r^{2}d\Omega^{2}.$ (111)
To examine the asymptotic form of $g_{rr}(r)$ for large $r$, we set $g_{rr}(r)=B_{0}r^{k}$, where $B_{0}$ is positive because of the condition (18), and study the Einstein tensors:
$-G^{t}{}_{t}(r)$ $=\frac{1}{r^{2}}+\frac{k-1}{B_{0}}r^{-2-k},$ (112)
$G^{r}{}_{r}(r)$ $=-\frac{1}{r^{2}}+\frac{3-k}{B_{0}}r^{-2-k},$ (113)
$G^{\theta}{}_{\theta}(r)$ $=\frac{k^{2}-4k+2}{2B_{0}}r^{-2-k}.$ (114)
As done in Sec.IV.2, we employ consistency with local thermodynamics (i.e. positivity of energy density and pressures). If $k<0$, $-G^{t}{}_{t}(r)$ would be negative for large $r$, meaning a negative energy density. Therefore, we get a condition $k\geq 0$. In order for the radial pressure to be positive for large $r$, we must have $r^{-2-k}\geq r^{-2}$ and $03-k>$ 0, leading to $k\leq 0$. Thus, we conclude $k=0$.
Then, the metric (111) reduces to
$ds^{2}=-\frac{C}{B_{0}}r^{2}dt^{2}+B_{0}dr^{2}+r^{2}d\Omega^{2},$ (115)
and the Einstein tensors become
$-G^{t}{}_{t}(r)=\frac{B_{0}-1}{B_{0}}r^{-2},~{}G^{r}{}_{r}(r)=\frac{3-B_{0}}{B_{0}}r^{-2},~{}G^{\theta}{}_{\theta}(r)=\frac{1}{B_{0}}r^{-2}.$ (116)
For these to be positive, we must have
$31<B_{0}<$ 3 . (117)
Then, the energy distribution is given from (13) by
$a(r)=\frac{B_{0}-1}{B_{0}}r<\frac{2}{3}r,$ (118)
which is smaller than the maximum one (24) (and Buchdahl’s limit, $\frac{8}{9}r$ Buchdahl ). Through the right hand side of (21), therefore, the entropy of the configuration (115) is $\sim\frac{B_{0}-1}{B_{0}}\frac{R^{2}}{l_{p}^{2}}<\frac{2}{3}\frac{R^{2}}{l_{p}^{2}}$. This cannot exceed $\sim\frac{R^{2}}{l_{p}^{2}}$ (25), the value estimated for the gravity condensate in the same way. Thus, we conclude that Tolman’s law does not maximize entropy for a given surface area.
Finally, we discuss an interesting point emerging as a byproduct. For $B_{0}=2$, (116) gives $-G^{t}{}_{t}=G^{r}{}_{r}=G^{\theta}{}_{\theta}=\frac{1}{2r^{2}}$. This represents Zel’dovich’s causal-limit fluid, where $\rho=p$ holds for $\rho=\langle-T^{t}{}_{t}\rangle,~{}p=\langle T^{r}{}_{r}\rangle=\langle T^{\theta}{}_{\theta}\rangle$ Zeldovich . The entropy $S\propto\frac{R^{2}}{l_{p}^{2}}$ can be obtained directly by applying the Gibbs relation $T_{loc}s_{3d}=\rho+p$, Tolman’s law $T_{loc}(r)=\frac{T_{0}}{\sqrt{-g_{tt}(r)}}$ and the metric (115) to the formula (2). This is a result supporting the consistency of our typicality argument. Also, the area-scaling entropy is consistent with Ref.Banks . It would be interesting to study the relation between their derivation and ours.